### v4

# Loss-Given-Default Modeling by post-last passage time process

Abstract.
This study proposes a stochastic model for loss-given-default (LGD) which provides the LGD distribution based on credit market and company-specific financial conditions. The model utilizes last passage time of a linear diffusion (representing firm value) to a certain threshold point, after which default occurs as a surprising event. By treating the post-last passage time process in a continuum of the original process, we are able to use firm-value approach before and intensity-based approach after the last passage time, leading to a hybrid model. Under minimal and standard assumptions, we obtain the distributions of default time and LGD explicitly. We provide a computationally simple estimation procedure and real-world examples of estimated LGD distribution implied in CDS market.
1Phone: +81-75-753-3430. 2Phone: +81-11-706-3174.
This version: November 2, 2025. An earlier version of this paper has been circulated under the title “A New Approach to Estimating Loss-Given-Default Distribution” or “Post-last exit time process and its application to loss-given-default distribution”. The first author is in part supported by Grant-in-Aid for Scientific Research (C) No.23K01467, Japan Society for the Promotion of Science. The second author is in part supported by JSPS KAKENHI Grant-in-Aid for Early-Career Scientists No.23K12501.
Keywords: loss given default distribution; credit risk; asset-to-debt ratio; linear diffusion; last passage time
Mathematics Subject Classification (2010): 60J60; 60J70
JEL classification: G32; C63
1. Introduction
This study is novel in that we have no assumptions on types of distribution or dynamics of loss-given-default (hereafter, LGD); yet, we are able to estimate the LGD distribution density, not merely an expected value. Our LGD distribution is implied in the CDS (credit default swap) market. As Gambetti et al. (2019) pointed out, general economic uncertainty has a considerable impact on company-specific LGD distribution. Considering this economic uncertainty is reflected in the price formation of the CDS market, a model that incorporates the market information should be of use and importance. Furthermore, an estimate of whole distribution density is appreciated from banks’ and investors’ credit risk management point of view.
There exists a large body of literature regarding the LGD distribution because the information about this distribution is important for credit risk management. One example of its use is the estimation of loan-loss reserves in the banking industry. The LGD distribution should reflect credit market conditions as well as company-specific financial conditions. However, as Doshi et al. (2018) and many other studies point out, it is a standard practice to assume, irrespective of company names, a constant recovery rate (approximately 40%) of the debt upon default. This $60\%$ loss rate is arbitrarily determined, possibly from the empirical distribution. Since the market data is usually quoted based on the predetermined constant loss rate (60%) for all firms, it is hard to estimate the LGD distribution by just observing the market information such as CDS spreads, default time distribution, and default probability. In contrast to the standard assumption of the constant loss rate, the empirical literature has documented the evidence of time-varying realized loss rates. Hence we cannot overemphasize the effectiveness of a model that provides the LGD distribution implied in the credit market and company-specific financial conditions.
This subject has long been an open issue. As the empirical literature has documented the evidence of time-varying realized loss rates, it is important to build a stochastic model for LGD. Gambetti et al. (2019) analyzes determinants of recovery rate distributions and finds economic uncertainty to be the most important systematic determinant of the mean and dispersion of the recovery rate distribution. For their analysis, the authors use post-default bond prices of 1831 American corporate defaults during 1990-2013. Even though recovery rates vary with a firm’s idiosyncratic factors, Gambetti et al. (2019) states that the impact of systematic factors related to economic cycles should not be underestimated. For the summary of studies related to cross-sectional and time variation of recovery rates, we refer the reader to Gambetti et al. (2019).
Altman et al. (2004) provides a detailed review of how recovery rate and its correlation with default probability had been treated in credit risk models. The authors also discuss the importance of modeling the correlation between recovery rate and default probability. Altman et al. (2005) analyzes and measures the relationship between default and recovery rates of corporate bonds over the period of 1982-2002. They confirm that default rate is a strong indicator of average recovery rate among corporate bonds. Acharya et al. (2007) analyzes data of defaulted firms in the U.S. during 1982-1999 and finds that the recovery rate is significantly lower when the industry of the defaulted firm is in distress. The authors discover that industry conditions at the time of default are robust and important determinants of recovery rates. Their results suggest that recovery rates are lower during industry distress not only because of the decreased worth of a firm’s assets but also because of the financial constraints that other firms in the industry face. The latter reason is based on the idea that the prices, at which the assets of the defaulted firm can be sold, depend on the financial condition of other firms in the industry.
Doshi et al. (2018) uses information extracted from senior and subordinate credit default swaps to identify risk-neutral stochastic recovery rate dynamics of credit spreads and studies the term structure of expected recovery. Their study is related to Schläfer and Uhrig-Homburg (2014) which also uses the fact that debt instruments of different seniority face the same default risk but have different recovery rates given default. Doshi et al. (2018) uses 5-factor intensity-based model for CDS contracts allowing stochastic dynamics of LGD. The authors allow firm-specific factors to influence the stochastic recovery rate. Their empirical analysis of 46 firms through the time period of 2001-2012 shows that the recovery rate is time-varying and the term structure of expected recovery is on average downward-sloping. They also find that industry characteristics have significant impact on CDS-implied recovery rates; however, they do not find the evidence that firms’ credit ratings explain the cross-sectional differences in recovery rates. For the summary of the literature related to the relationship between default rates and realized recovery rates, refer to Doshi et al. (2018).
Yamashita and Yoshiba (2013) is an example of a study that uses stochastic collateral value process to incorporate stochastic recovery rate into the model which assumes that a constant portion of the collateral value is recovered upon default. Yamashita and Yoshiba (2013) uses a quadratic Gaussian process for the default intensity and discount interest rate and derives an analytical solution for the expected loss and higher moments of the discounted loss distribution for a collateralized loan. The authors assume that the default intensity, discount interest rate, and collateral value are correlated through Brownian motions driving Gaussian state variables.
Finally, Cohen and Costanzino (2017) is an example of a study that extends a structural credit risk model and incorporates stochastic dynamics of the recovery rate. In contrast to our approach which does not introduce an additional recovery risk factor and is based solely on the dynamics of the asset-to-debt ratio process, Cohen and Costanzino (2017) models asset and recovery processes separately as correlated geometric Brownian motions. In their model, the asset risk driver serves as a default trigger and the recovery risk driver determines the amount recovered upon default. The authors explicitly compute the prices of bonds and CDS under this framework. See also Kijima et al. (2009).
We take a completely different approach to modeling LGD distribution: using the post-last passage time process. Let $X=\{X_{t},t\geq 0\}$ be a linear diffusion. We study the process after its last passage time from a certain regular point: $(X_{t})_{t>\lambda_{\alpha}}$ where $\lambda_{\alpha}$ is the last passage time from state $\alpha$. We call this $(X_{t})_{t>\lambda_{\alpha}}$ the post-last passage time process. The post-last passage time process in Section 3.2 is, to our knowledge, first studied by Meyer et al. (1972) where the process $(X_{\lambda_{\alpha}+t},\mathcal{F}_{\lambda_{\alpha}+t})$ is proved to be a strong Markov process and its transition semigroup is identified. While it is treated as a newborn process at time $\lambda_{\alpha}$ in Meyer et al. (1972), we treat the post-last passage time process in a continuum of the original process $X$: we look at the entire path, a contributing factor for the modeling of LGD distribution.
Remark 1.1 (Economic interpretation of last passage time).
The model is discussed in Section 2 and an empirical study is performed in Section 5. The essential rationale of using last passage time is as follows. While the firm-value approach for credit risk analysis proposed in Merton (1974) is a fundamental method (as is the intensity-based approach) in estimating the default probability, one has some difficulties in calculating the LGD under this approach. For example, if the firm’s default is defined as the first hitting time of the firm value to a certain point, say $c$, then the value upon default is necessarily equal to the deterministic number $c$. To overcome this discrepancy, we consider the last passage time of the firm value (measured by its asset-to-debt ratio) to another threshold level $\alpha>c$. First, the fact that the firm value will not recover to level $\alpha$ after time $\lambda_{\alpha}$ is significant since it implies that the company is not creditworthy after $\lambda_{\alpha}$. We shall make use of this observation. Moreover, this approach has a useful feature. Since the last passage time is not a stopping time, it is convenient when modeling the reality: during the period while the observer does not know whether the firm value would recover to $\alpha$, a default occurs as a surprise. Moreover, the post-last passage time process well captures this reality and is suitable for estimating LGD distribution. In fact, we compute LGD distribution through the intensity-based approach applied to the post-last passage time process. In this modeling, due to the mathematical contribution (in Section 3.2) which allows us to look at the entire path of $X$ as mentioned above, we can deal with the firm value process consistently and continuously: the firm-value approach before time $\lambda_{\alpha}$ and intensity-based approach after $\lambda_{\alpha}$. Another advantage of the modeling is that we do not need to detect or predict when the last passage time occurs in the estimation of LGD distribution.
In Section 4, with minimal and standard assumptions, the model to be presented allows us to obtain the LGD distribution implied in the current CDS market: we compute (1) the model-driven LGD distribution and (2) CDS spreads implied by the estimated LGD distribution. The results are the explicit formula for the distribution of default time in Proposition 4.1 and that of LGD in Corollary 4.2. We also include a procedure to make sure whether the estimated LGD reflects the market data. When checking the consistency of the model-implied CDS spreads with the market quotes, we compare “spreads per 1% loss given default” because the market CDS spreads are quoted under the assumption that the LGD is fixed at 60% for all firms. In the end, we emphasize that one should not confuse our loss right upon the default with the final loss after liquidation and/or bankruptcy procedure. For our analysis, the risk-free rates were obtained from the website of the Federal Reserve Bank of St.Louis. The remaining data was obtained from LSEG Workspace.
Section 5 is devoted to empirical study, where we compute three probability densities; in addition to the LGD distribution (Figure 1) of eight companies, we show the distribution of the last passage time $\lambda_{\alpha}$ (Figure 2) and the occupancy time above level $\alpha$ before the process is killed at some levels of $c<\alpha$ (Figure 3). The former is based on the study in Salminen (1984) and Egami and Kevkhishvili (2020), and the latter is a new result derived in Proposition 3.1. All three distributions are, we believe, important and useful for risk management.
The last passage time is an important subject in the probability literature due to being closely related to the concepts of transience/recurrence, Doob’s $h$-transform, time-reversed processes, and the Martin boundary theory. For these interrelated topics, we refer the reader to Doob (1957), Nagasawa (1964), Kunita and Watanabe (1966), Salminen (1984), Rogers and Williams (1994a), Borodin and Salminen (2002), Chung and Walsh (2004), Revuz and Yor (2005) as well as the literature referenced therein. While the last passage time is not a stopping time, it has been utilized widely in financial applications as discussed in Nikeghbali and Platen (2013). These applications include credit risk analysis (Egami and Kevkhishvili, 2020), valuation of defaultable claims (Elliott et al. (2000), Jeanblanc and Rutkowski (2000), Coculescu and Nikeghbali (2012), Jeanblanc et al. (2009)), option valuation (Profeta et al. (2010), Cheridito et al. (2012)), and insider trading (Imkeller, 2002).
2. Loss-given-default Modeling
2.1. Basic idea
The goal is to obtain the LGD distribution implied in the current CDS market and company-specific financial conditions. More specifically, we derive the distribution of the ratio of the firm’s total assets over its total debt upon default. We shall hereafter refer to the dynamics of this ratio as the asset-to-debt ratio process, denoted by $Y$. In our framework, we are focusing on the loss rate of the debt of the whole company, not each individual debt obligation. Based on this corporate-level loss distribution, credit managers can calculate loss rates for individual debt obligations having distinct characteristics; for example, senior or subordinated, and with or without collateral. This is convenient for empirical study because the spreads in the CDS market are quoted in terms of the risk of the corporate name, not that of its specific debt obligation.
Fix a probability space $(\Omega,\mathcal{H},\mathbb{P})$ and a filtration $\mathcal{F}=(\mathcal{F}_{t})_{t\geq 0}$. Let $X=\{\omega(t),t\geq 0;\mathbb{P}_{x}\}$ be a linear regular canonical diffusion starting at $x\in\mathbb{R}$: that is, $\omega(t)=X_{t}(\omega),t\geq 0$. Its state space is given by $\mathcal{I}=(\ell,r)\subset\mathbb{R}$. Let the last passage time from a regular point $\alpha\in\mathcal{I}$ be denoted by
$\lambda_{\alpha}:=\sup\{t:\omega(t)=\alpha\}$
with $\sup\emptyset=0$. The last passage time is an example of co-optional time satisfying
$\lambda_{\alpha}\circ\theta_{t}=(\lambda_{\alpha}-t)^{+},\quad t\geq 0,$
where $\{\theta_{t},t\geq 0\}$ is the shift operator in the following manner. For each $t$, $\theta_{t}$ maps $\Omega$ into $\Omega$ such that
$\forall t:(X_{s}\circ\theta_{t})(\omega)=X_{s}(\theta_{t}\omega)=X_{t+s}(\omega).$
In the firm-value (or Merton (1974)) model of credit risk, insolvency occurs upon the asset-to-debt ratio being equal to some level $c$. In this case, if we set $\alpha>c$, then there must exist the last passage time $\lambda_{\alpha}$ recorded by the asset-to-debt ratio process (to be defined in the next Section 2.3). The idea is summarized as follows:
  1. (1)
    We choose $\alpha$ so that the default (insolvency to be exact) probability implied by our model is consistent with that implied by the CDS market. Hence the level $\alpha$ is calibrated to the market: there is no arbitrariness in the selection of $\alpha$.
  2. (2)
    Since $\lambda_{\alpha}$ is not a stopping time, it is appropriate to model the interval, denoted by $\tau$, between $\lambda_{\alpha}$ and default time $\xi$ by the intensity-based model:
    $\xi(\omega):=\lambda_{\alpha}(\omega)+\tau\circ\theta(\lambda_{\alpha}(\omega)).$
    In fact, this well explains the reality that a default occurs surprisingly during the period when non-insider investors do not have enough information about credit quality. The $\lambda_{\alpha}$, not a stopping time, corresponds to the time when creditworthiness becomes hard to grasp, and $\xi$ occurs as a surprise after non-predicting $\tau$ has elapsed.
  3. (3)
    The company’s asset-to-debt ratio at $\xi$ is then computed as the distribution of the value of the asset-to-debt ratio process at time $\xi$, from which the LGD distribution at the corporate level is obtained. For this purpose, we use results from Section 3.2.
We explicitly derive the distribution of default time $\xi$ (Proposition 4.1) and LGD (Corollary 4.2). These are computationally simple formulas. Furthermore, the model check proposed in Section 4.3 inspects whether the model-implied default time and LGD distribution are consistent with the credit market and company-specific financial conditions. Specifically, we check that the CDS spread implied from the estimated LGD distribution is consistent with the quoted spread in the market.
Let us emphasize that we treat the entire process in a continuum starting at time zero, rather than treating the process after its passage time as a newborn. This is important since for practical purposes we wish to observe the process through its lifetime (from time zero to the killing). It is particularly true when we look at a company’s creditworthiness and estimate its LGD distribution.
2.2. Mathematical Setting
Let us return to the general mathematical setup at the beginning of this section for a general diffusion $X$. We refer the reader to Chapter II in Borodin and Salminen (2002) for the basic facts regarding linear diffusions. The scale function and the speed measure of $X$ are denoted by $s(\cdot)$ and $m(\cdot)$, respectively. We assume that $s$ and $m$ are absolutely continuous with respect to the Lebesgue measure and have smooth derivatives:
$m({\rm d}x)=m(x){\rm d}x\quad\text{and}\quad s(x)=\int^{x}s^{\prime}(y){\rm d}y,$
where $m$ and $s^{\prime}$ are continuous and positive. If $X$ hits $\ell$ or $r$, it is killed and transferred immediately to the cemetery $\Delta\notin\mathcal{I}$. We assume that $X$ is transient. In this paper, we call a diffusion transient (following Salminen (1984)) if
$\forall x\in\mathcal{I},\;\text{$A$ is compact in $\mathcal{B}(\mathcal{I})$: }\mathbb{P}_{x}(\lambda_{A}<\infty)=1\;\textnormal{with}\;\lambda_{A}=\sup\{t:\omega(t)\in A\}.$
The transience in the above sense is equivalent to one or both of the boundaries being attracting; that is, $s(\ell)>$-$\infty$ and/or $s(r)<+\infty$. To obtain concrete results, we set a specific assumption:
Assumption 1.
$s(\ell)>-\infty\quad\text{and}\quad s(r)=+\infty.$
Then, it holds that
$\mathbb{P}_{x}\left(\lim_{t\to\xi}\omega(t)=\ell\right)=1,\quad\forall x\in\mathcal{I}.$
Hence the process is killed at the lower boundary $\ell$, which is consistent with our asset-to-debt ratio process.
The infinitesimal drift and diffusion parameters are given by $\mu(\cdot)$ and $\sigma(\cdot)$, respectively. We let $\mathcal{G}$ denote the second-order differential operator
$\mathcal{G}f(x)=\frac{1}{2}\sigma^{2}(x)f^{\prime\prime}(x)+\mu(x)f^{\prime}(x),\quad x\in\mathcal{I}.$
For each $t\geq 0$, the transition function is given by $P_{t}:\mathcal{I}\times\mathcal{B}(\mathcal{I})\mapsto[0,1]$ such that for all $t,s\geq 0$
$\mathbb{P}\left(X_{t+s}\in A\mid\mathcal{F}_{s}\right)=P_{t}(X_{s},A),\quad\forall A\in\mathcal{B}(\mathcal{I})\quad\mathbb{P}-a.s.$
For each $t>0$ and $x\in\mathcal{I}$, $P_{t}(x,\cdot):A\mapsto P_{t}(x,A)$ is absolutely continuous with respect to the speed measure $m$:
$P_{t}(x,A)=\int_{A}p(t;x,y)m({\rm d}y),\quad A\in\mathcal{B}(\mathcal{I}).$
The transition density $p$ may be taken to be jointly continuous in all variables and symmetric such that $p(t;x,y)=p(t;y,x)$. Note that the Green function $G(y,z)$ for $y,z\in\mathcal{I}$ has a representation in terms of the scale function:
$G(y,z):=\int_{0}^{\infty}p(t;y,z){\rm d}t=\begin{cases}\lim\limits_{a\downarrow\ell,b\uparrow r}\frac{(s(y)-s(a))(s(b)-s(z))}{s(b)-s(a)},\quad\ell<y\leq z<r,\\ \lim\limits_{a\downarrow\ell,b\uparrow r}\frac{(s(z)-s(a))(s(b)-s(y))}{s(b)-s(a)},\quad\ell<z\leq y<r.\end{cases}$ (2.1)
The Laplace transform of the hitting time $H_{z}:=\inf\{t\geq 0:X_{t}=z\}$ for $z\in\mathcal{I}$ is given by
$\mathbb{E}_{x}\left[e^{-qH_{z}}\right]=\begin{cases}\frac{\phi_{q}(x)}{\phi_{q}(z)},\quad x\geq z,\\ \frac{\psi_{q}(x)}{\psi_{q}(z)},\quad x\leq z,\end{cases}$ (2.2)
where the continuous positive functions $\psi_{q}$ and $\phi_{q}$ represent linearly independent solutions of the ODE $\mathcal{G}f=qf$ with $q>0$. Note that $\psi_{q}$ is increasing while $\phi_{q}$ is decreasing. These functions are unique up to a multiplicative constant, once the boundary conditions at $\ell$ and $r$ are specified. Finally, the q-Green function is defined as
$G_{q}(x,y):=\begin{cases}\frac{\psi_{q}(y)\phi_{q}(x)}{w_{q}},\quad x\geq y,\\ \frac{\psi_{q}(x)\phi_{q}(y)}{w_{q}},\quad x\leq y,\end{cases}$ (2.3)
with the Wronskian $w_{q}:=\psi_{q}^{+}(x)\phi_{q}(x)-\psi_{q}(x)\phi_{q}^{+}(x)=\psi_{q}^{-}(x)\phi_{q}(x)-\psi_{q}(x)\phi_{q}^{-}(x)$. It holds that $G_{q}(x,y)=\int_{0}^{\infty}e^{-qt}p(t;x,y){\rm d}t$ for $x,y\in\mathcal{I}$.
2.3. Asset-to-debt ratio process
Before providing mathematical results necessary for our analysis in Section 3.2 below, we define the asset-to-debt ratio process which plays the central role in our model. Let us assume that the firm value (market value of total assets) $V$ follows a geometric Brownian motion:
${\rm d}V_{t}=\mu V_{t}{\rm d}t+\sigma V_{t}{\rm d}W_{t}$ (2.4)
with $\mu\in\mathbb{R}$ and $\sigma>0$. The debt process is given by
$B_{t}=B_{0}e^{rt},\quad B_{0}\in\mathcal{F}_{0},\quad t\geq 0,$
where $r>0$ denotes a constant rate. As announced above, we define the asset-to-debt ratio process $Y=V/B$ on $(0,+\infty)$:
$Y_{t}=Y_{0}e^{(\mu-\frac{1}{2}\sigma^{2}-r)t+\sigma W_{t}}\quad\text{and}\quad\ln\left(Y_{t}\right)=\ln\left(Y_{0}\right)+\left(\mu-\frac{1}{2}\sigma^{2}-r\right)t+\sigma W_{t}.$
The parameters $\mu$, $\sigma$ are such that the left boundary 0 is attracting and the right boundary $\infty$ is non-attracting, which ensures that the default occurs eventually with probability one. This complies with Assumption 1. In particular, we assume that $\mu-\frac{1}{2}\sigma^{2}-r<0$. By fixing a certain level $\alpha\in\mathbb{R}_{+}$, we define
$\lambda_{\alpha}:=\sup\{t\geq 0:Y_{t}=\alpha\}.$
We have the normalized process, with state space $\mathcal{I}=(\ell,r)=(-\infty,+\infty)$,
$\frac{1}{\sigma}{\rm d}\ln\left(Y_{t}\right)=\frac{\mu-\frac{1}{2}\sigma^{2}-r}{\sigma}{\rm d}t+{\rm d}W_{t}=M{\rm d}t+{\rm d}W_{t}$ (2.5)
by setting $M:=\frac{\mu-\frac{1}{2}\sigma^{2}-r}{\sigma}$. Now let us introduce the post-last passage time process
$Z_{t}:=\left(\frac{1}{\sigma}\ln(Y_{t})\right)\circ\theta(\lambda_{\alpha}),\quad t>0.$
We need first to identify the dynamics of the process $Z$ to estimate its value at the default time $\xi$, which we shall undertake in Section 3.2 after first providing new mathematical results in the next section.
3. Two mathematical results concerning last passage time
In this section, let us insert two mathematical results that apply to general transient linear diffusions. Both are, we believe, interesting in their own rights (since the results are general) and useful for the purpose of credit risk modeling. Due to the fact that our proposed modeling handles a whole lifetime of diffusion (the asset-to-debt ratio process), pre- and post- last passage time state can be consistently analyzed with these mathematical results.
3.1. Occupancy time above $\alpha$ before the last passage time
The first result is concerned with the occupation time above a certain level, say $\alpha\in\mathcal{I}=(\ell,r)$, before transient diffusion is killed. There is a well-known example in the classic textbook of Karlin and Taylor (1981) (see its Section 8-B, Chapter 15) where a diffusion is killed at its first passage to the left boundary and the time change by the occupation time of the positive axis is considered. The last passage time $\lambda_{\alpha}$ represents the final instance when the process is in the area above level $\alpha$. We provide the distribution, in terms of its Laplace transform, of the occupancy time above $\alpha$ for a general transient diffusion (A recurrent diffusion can be made transient if we set some finite state as a killing boundary). To our knowledge, this distribution is not available in the literature.
In the context of credit risk, this translates into how long the company’s asset-to-debt ratio stays above level $\alpha$. We can compute the distribution just by using estimated parameters in (2.5). This information should be of value to the management, banks, and investors. A ready-to-go practice is to set the killing boundary $\ell=0.4$ for the process $\frac{1}{\sigma}{\rm d}\ln(Y_{t})$, assuming that its LGD is around the market convention of $60\%$.
Fix some $\alpha\in\mathcal{I}$ as usual. Consider an occupation time of the region above $\alpha$
$\Gamma_{+}(t):=\int_{0}^{t}\mathbf{1}_{\{X_{s}\geq\alpha\}}{\rm d}s,\quad t\geq 0,$
together with its right inverse:
$\Gamma_{+}^{-1}(t):=\inf\{s:\Gamma_{+}(s)>t\}.$
Now let us define $\lambda^{A}_{\alpha}:=\Gamma_{+}(\lambda_{\alpha})$, which is the occupation time of $X$ above level $\alpha$ up to the last passage time $\lambda_{\alpha}$. Under Assumption 1, it is almost sure finite. We will use the following time-changed process:
$\hat{X}^{A}(t):=X(\Gamma_{+}^{-1}(t)).$
This process $\hat{X}^{A}$ is illustrated as FIG. 4 in the aforementioned example of Karlin and Taylor (1981) and the letter $\zeta$ there should read $\lambda^{A}_{\alpha}$ in the current notation.
Let us introduce another diffusion $X^{A}$ on $\alpha,r)$ for which $\alpha$ is a reflecting boundary. The process $X^{A}$ has the same speed measure and scale function as $\hat{X}^{A}$ (hence the same as $X$ as shown in Dynkin (1965, Theorem 10.12)) but its killing measure is zero. To be precise, we denote
$\hat{X}^{A}(t)=\begin{cases}X^{A}(t),\quad 0\leq t<\lambda_{\alpha}^{A}\\ \Delta,\quad t\geq\lambda_{\alpha}^{A}\end{cases}$
to distinguish the killed process $\hat{X}^{A}$ from $X^{A}$.
Before we proceed further, we need to introduce the local time at $\alpha$ for $\hat{X}^{A}$ by denoting
$\hat{L}^{A}(t):=\hat{L}^{A}(t,\alpha)=L(\Gamma^{-1}_{+}(t),\alpha)$
where $L(\cdot,\alpha)$ is the local time of $X$ at $\alpha$ which we denote by $L(\cdot)$ for brevity. We note that $X^{A}(t)=\hat{X}^{A}(t)$ and $L^{A}(t)=\hat{L}^{A}(t)$ on $0,\lambda_{\alpha}^{A}]$ where $L^{A}(t)$ is the local time at $\alpha$ for $X^{A}$. Due to this fact, we can deal with $L^{A}(\cdot)$ rather than $\hat{L}^{A}(\cdot)$ in the proof of the next proposition.
Let us also define the inverse local time process $\rho^{A}(s):=\inf\{t:L^{A}(t)>s\}$ for $X^{A}$ (see (Getoor, 1979, Section 7)) and also the q-Green function of $X^{A}$, denoted by $G_{q}^{A}$. The latter is, after some computation, $G_{q}^{A}(\alpha,\alpha)=\frac{\phi_{q}(\alpha)}{-\phi_{q}^{+}(\alpha)}$ where $\phi_{q}$ is in (2.2) and $\phi_{q}^{+}$ is the right derivative with respect to the scale function $s$.
Proposition 3.1.
Suppose that the state space of a general transient linear diffusion $X$ is $\mathcal{I}=(\ell,r)$. Under Assumption 1, the Laplace transform of $\lambda^{A}_{\alpha}$, the occupation time above level $\alpha$ before the last passage time $\lambda_{\alpha}$, is
$\mathbb{E}_{x}\left[e^{-q\lambda^{A}_{\alpha}}\right]=\frac{G^{A}_{q}(x,\alpha)}{G^{A}_{q}(\alpha,\alpha)+G(\alpha,\alpha)},\quad q>0,\quad x\geq\alpha,$ (3.1)
where $G(\alpha,\alpha)=s(\alpha)-s(\ell)$ (see (2.1)). Note that $\ell$ can be a finite ($\alpha>\ell$) number or negative infinity.
Proof.
Let $\mathbf{e_{q}}$ be an exponential random variable, independent of $X$, with rate $q>0$. By the independence of $X$ and $\mathbf{e_{q}}$, we have
$\mathbb{E}_{\alpha}\left[e^{-q\lambda^{A}_{\alpha}}\right]=\mathbb{P}_{\alpha}(\lambda^{A}_{\alpha}<\mathbf{e_{q}})=\mathbb{P}_{\alpha}(L^{A}(\lambda^{A}_{\alpha})<L^{A}(\mathbf{e_{q}})).$
Since $\lambda^{A}_{\alpha}:=\Gamma_{+}(\lambda_{\alpha})$ and $L^{A}(\lambda^{A}_{\alpha})=L(\Gamma^{-1}_{+}(\lambda^{A}_{\alpha}))$, we have
$L^{A}(\lambda^{A}_{\alpha})=L(\Gamma^{-1}_{+}(\Gamma_{+}(\lambda_{\alpha}))=\lim_{t\rightarrow\lambda_{\alpha}-}L(t)=L(\lambda_{\alpha}).$
By the argument in Section A.1, $L(\lambda_{\alpha})<\infty$ is the local time when the first excursion (from $\alpha$) of infinite length occurs and is exponentially distributed, as a Poisson arrival time, with rate $\frac{1}{G(\alpha,\alpha)}$. In other words, corresponding to local time $L(\lambda_{\alpha})$, we have a point (of the Poisson point process) representing that length of excursion.
On the other hand, it is known that the random variable $L^{A}(\mathbf{e_{q}})$ is exponentially distributed and
$\mathbb{P}_{\alpha}(L^{A}(\mathbf{e_{q}})>s)=\mathbb{P}_{\alpha}(\rho^{A}(s)<\mathbf{e_{q}})=\mathbb{E}_{\alpha}\left[e^{-q\rho^{A}(s)}\right]=\exp\left(-\frac{s}{G^{A}_{q}(\alpha,\alpha)}\right).$
The random variable $L(\lambda_{\alpha})$ is independent of any occasion that corresponds to a certain length of excursion because of the Markov property employed at each time an excursion begins. In particular, it is independent of $L^{A}(\mathbf{e_{q}})$. Therefore, we can compute
$\mathbb{P}_{\alpha}(L^{A}(\lambda^{A}_{\alpha})<L^{A}(\mathbf{e_{q}}))=\frac{\frac{1}{G(\alpha,\alpha)}}{\frac{1}{G(\alpha,\alpha)}+\frac{1}{G^{A}_{q}(\alpha,\alpha)}},$
which provides (3.1) with $x=\alpha$. For a general $x>\alpha$, by the strong Markov property at $H_{\alpha}$, we compute
$\mathbb{E}_{x}[e^{-q\lambda_{\alpha}}]=\mathbb{E}_{x}[e^{-q(H_{\alpha}+\lambda_{\alpha}\circ\theta_{H_{\alpha}})}]=\mathbb{E}_{x}[e^{-qH_{\alpha}}]\cdot\mathbb{E}_{\alpha}[e^{-q\lambda_{\alpha}}]$
and use $\mathbb{E}_{x}[e^{-qH_{\alpha}}]=\frac{\phi_{q}(x)}{\phi_{q}(\alpha)}$ (see (2.2)) and (2.3) to obtain (3.1). Note that $\phi_{q}^{A}(x)=\phi_{q}(x)$ on $\alpha,r)$ where $\phi_{q}^{A}$ denotes the decreasing solution of $\mathcal{G}^{A}f=qf$ with the generator $\mathcal{G}^{A}$ of $X^{A}$. ∎
For a comment on computations of $G_{q}^{A}(\cdot,\cdot)$ in various cases, see Appendix A.2.
3.2. Post-last passage time process
In this subsection, we shall identify a general form of the transition density of post-last passage time process in the class of general transient linear diffusions.
We are interested in the distribution of a general transient diffusion $X$ after the last passage time from $\alpha$; i.e., $\mathbb{P}_{x}(X_{t}\in{\rm d}y,\lambda_{\alpha}<t)$. Recall that for transient diffusion $\lambda_{\alpha}<\infty$ a.s. Let us define
$h_{\alpha}(x):=\mathbb{P}_{x}(\lambda_{\alpha}=0)$ (3.2)
and a transform of the transition density of $X$ by
$p^{h_{\alpha}}(t;x,y):=\frac{h_{\alpha}(y)}{h_{\alpha}(x)}p(t;x,y),\quad h_{\alpha}(x)\neq 0$ (3.3)
with respect to the speed measure $m({\rm d}y)$.
Proposition 3.2.
Suppose that the state space of a general transient linear diffusion $X$ is $\mathcal{I}=(\ell,r)$. Under Assumption 1, the transition density (with respect to the speed measure $m({\rm d}y)$) of the post-last passage time process $(X_{t})_{t>\lambda_{\alpha}}$ is given by (3.3). More precisely,
$p^{h_{\alpha}}(t-u;z,y)=\frac{h_{\alpha}(y)}{h_{\alpha}(z)}p(t-u;z,y),\qquad y<z<\alpha,\quad u\in(\lambda_{\alpha},t)$ (3.4)
for $u$ arbitrarily close to $\lambda_{\alpha}$: $z$ is arbitrarily close to $\alpha$.
Proof.
[…]
$\mathbb{P}_{x}(X_{t}\in{\rm d}y,\lambda_{\alpha}<t)$ $=\mathbb{P}_{x}(\lambda_{\alpha}<t\mid X_{t}\in{\rm d}y)\cdot\mathbb{P}_{x}(X_{t}\in{\rm d}y)=\mathbb{P}_{x}(\lambda_{\alpha}\circ\theta_{t}=0|X_{t}\in{\rm d}y)\cdot\mathbb{P}_{x}(X_{t}\in{\rm d}y,H_{z}<t)$
$=\mathbb{P}_{y}(\lambda_{\alpha}=0)\cdot\mathbb{E}_{x}\left[\mathbb{E}[\mbox{1}\hskip-2.5pt\mbox{l}_{\{X_{t}\in{\rm d}y\}}\cdot\mbox{1}\hskip-2.5pt\mbox{l}_{\{H_{z}<t\}}\mid\mathcal{F}_{H_{z}}]\right]$
$=h_{\alpha}(y)\cdot\mathbb{E}_{x}\left[p(t-H_{z};X_{H_{z}},y)m({\rm d}y)\cdot\mbox{1}\hskip-2.5pt\mbox{l}_{\{H_{z}<t\}}\right]$
$=h_{\alpha}(y)\cdot\mathbb{E}_{x}\left[\left(p(t-H_{z};X_{H_{z}},y)\mbox{1}\hskip-2.5pt\mbox{l}_{\{\lambda_{\alpha}<H_{z}<t\}}+p(t-H_{z};X_{H_{z}},y)\mbox{1}\hskip-2.5pt\mbox{l}_{\{H_{z}<\lambda_{\alpha}<t\}}\right)m({\rm d}y)\right]$
where we used Markov property in the second and strong Markov property together with (3.2) in the third line. Note that $\mathbb{E}_{x}\left[h_{\alpha}(y)\cdot p(t-H_{z};X_{H_{z}},y)\cdot\mbox{1}\hskip-2.5pt\mbox{l}_{\{\lambda_{\alpha}>t>H_{z}\}}\right]=\mathbb{E}_{x}\left[\mathbb{E}_{x}\left[\mathbb{P}_{y}(\lambda_{\alpha}=0)\cdot\mbox{1}\hskip-2.5pt\mbox{l}_{\{X_{t}\in{\rm d}y\}}\cdot\mbox{1}\hskip-2.5pt\mbox{l}_{\{H_{z}<t\}}\mid\mathcal{F}_{H_{z}}\right]\cdot\mbox{1}\hskip-2.5pt\mbox{l}_{\{\lambda_{\alpha}>t\}}\right]=0.$ Let us examine separately the two terms on the right-hand side. The first term implies that $H_{z}$ occurs after $\lambda_{\alpha}$. Hence we use the transform (3.3) to obtain
$\mathbb{E}_{x}\left[p(t-H_{z};X_{H_{z}},y)\mbox{1}\hskip-2.5pt\mbox{l}_{\{\lambda_{\alpha}<H_{z}<t\}}\right]$ $=\mathbb{E}_{x}\left[\frac{h_{\alpha}(X_{H_{z}})}{h_{\alpha}(y)}p^{h_{\alpha}}(t-H_{z};X_{H_{z}},y)\mbox{1}\hskip-2.5pt\mbox{l}_{\{\lambda_{\alpha}<H_{z}<t\}}\right]=\mathbb{E}_{x}\left[\frac{1}{h_{\alpha}(y)}p^{h_{\alpha}}(t-T_{z};X_{T_{z}},y)\mbox{1}\hskip-2.5pt\mbox{l}_{\{\lambda_{\alpha}<H_{z}<t\}}\right]$
where the last equality is due to $h_{\alpha}(X_{H_{z}})=\mathbb{P}_{X_{H_{z}}}(\lambda_{\alpha}=0)=1$ and $H_{z}=T_{z}$ on the set $\{\lambda_{\alpha}<H_{z}<t\}$.
To deal with the second term, we define, for an arbitrary random time $T<\infty$ a.s., a random measure $\tilde{p}(T;u,v)$ such that $\mathbb{E}_{x}[\tilde{p}(T;u,v)m({\rm d}v)]=\mathbb{P}_{u}(X_{T}\in{\rm d}v)$ for $u,v\in\mathcal{I}$. Note that $\tilde{p}(T_{z};x,v)m({\rm d}v)=0$ a.s. when $v\neq z$. This is because by Fubini’s theorem
$\mathbb{E}_{x}\left[\int_{\mathcal{I}}\tilde{p}(T_{z};x,v)m({\rm d}v)\right]$ $=\int_{\mathcal{I}}\mathbb{P}_{x}(X_{T_{z}}\in{\rm d}v)=\int_{\mathcal{I}}\int_{0}^{\infty}\mathbb{P}_{x}(X_{s}\in{\rm d}v,T_{z}\in{\rm d}s)$
$=\int_{0}^{\infty}\mathbb{P}_{x}(X_{s}\in{\rm d}z,T_{z}\in{\rm d}s)=\mathbb{P}_{x}(X_{T_{z}}\in{\rm d}z)=\mathbb{E}_{x}[\tilde{p}(T_{z};x,v)\delta_{z}({\rm d}v)]$
and $\mathbb{P}_{x}(X_{T_{z}}\in{\rm d}z)=1$ by the definition of $T_{z}$. Hence the random measure $\tilde{p}(T_{z};x,v)m({\rm d}v)$ has a point mass at $z$. With this preparation the second term becomes
$\mathbb{E}_{x}\left[p(t-H_{z};X_{H_{z}},y)\mbox{1}\hskip-2.5pt\mbox{l}_{\{H_{z}<\lambda_{\alpha}<t\}}\right]$ $=\mathbb{E}_{x}\left[\int_{\mathcal{I}}\tilde{p}(T_{z}-H_{z};z,u)\tilde{p}(t-T_{z};u,y)m({\rm d}u)\mbox{1}\hskip-2.5pt\mbox{l}_{\{H_{z}<\lambda_{\alpha}<T_{z}<t\}}\right]$
$=\mathbb{E}_{x}\left[p(t-T_{z};X_{T_{z}},y)\mbox{1}\hskip-2.5pt\mbox{l}_{\{H_{z}<\lambda_{\alpha}<T_{z}<t\}}\right]$
$=\mathbb{E}_{x}\left[\frac{h_{\alpha}(X_{T_{z}})}{h_{\alpha}(y)}p^{h_{\alpha}}(t-T_{z};X_{T_{z}},y)\mbox{1}\hskip-2.5pt\mbox{l}_{\{H_{z}<\lambda_{\alpha}<T_{z}<t\}}\right]$
$=\mathbb{E}_{x}\left[\frac{1}{h_{\alpha}(y)}p^{h_{\alpha}}(t-T_{z};X_{T_{z}},y)\mbox{1}\hskip-2.5pt\mbox{l}_{\{H_{z}<\lambda_{\alpha}<T_{z}<t\}}\right],$
where $h_{\alpha}(X_{T_{z}})=1$ is due to the definition of $T_{z}$. Note that $\mbox{1}\hskip-2.5pt\mbox{l}_{\{\lambda_{\alpha}<H_{z}<t\}}+\mbox{1}\hskip-2.5pt\mbox{l}_{\{H_{z}<\lambda_{\alpha}<T_{z}<t\}}=\mbox{1}\hskip-2.5pt\mbox{l}_{\{T_{z}<t\}}$ by the assumption that $T_{z}<t$ whenever $\lambda_{\alpha}<t$. Combine these results to obtain
$\mathbb{P}_{x}(X_{t}\in{\rm d}y,\lambda_{\alpha}<t)=\mathbb{E}_{x}\left[p^{h_{\alpha}}(t-T_{z};X_{T_{z}},y)\mbox{1}\hskip-2.5pt\mbox{l}_{\{T_{z}<t\}}m({\rm d}y)\right]=\int_{0}^{t}p^{h_{\alpha}}(t-s;z,y)\mathbb{P}_{x}(T_{z}\in{\rm d}s)m({\rm d}y).$
Since $z$ is arbitrary, we can make $z$ as close to $\alpha$ as we please. Therefore, for any time $\lambda_{\alpha}+u,u>0$, the post- last passage time process is governed by (3.4) for $u$ arbitrarily close to zero. ∎
Although $h_{\alpha}(x)$ is not an excessive function, the transform (3.4) has the same form as if it were an excessive function. (See the literature listed in Section 1 for Doob’s $h$-transform.) Therefore, we can compute the infinitesimal drift and diffusion parameters of the transform (3.3) in the same way as if it is an $h$-transform. But there is a caveat that $x\neq\alpha$.
Suppose that $X$ is a diffusion (does not have to be transient) and let $h$ be an excessive function. Let us denote the infinitesimal drift and diffusion parameters of the $h$-transform by $\mu^{h}(\cdot)$ and $\sigma^{h}(\cdot)$, respectively. The following result slightly generalizes the argument in Section 15.9 of Karlin and Taylor (1981).
Lemma 3.1.
Assume $h$ is an excessive function and its derivative $h^{\prime}(y)$ exists for $y\in\mathcal{I}$. The infinitesimal drift $\mu^{h}(y)$ and diffusion parameter $\sigma^{h}(y)$ of the $h$-transform of $Y$ are
$\mu^{h}(y)=\mu(y)+\frac{h^{\prime}(y)}{h(y)}\sigma^{2}(y)\quad\text{and}\quad\sigma^{h}(y)=\sigma(y).$
Proof.
See Appendix A.3. ∎
Turning to our case of $h_{\alpha}$,
$h_{\alpha}(x)=\mathbb{P}_{x}(\lambda_{\alpha}=0)=\frac{s(\alpha)-s(x)}{s(\alpha)-s(\ell)},\quad x\leq\alpha.$ (3.5)
We have $\frac{h_{\alpha}^{\prime}(x)}{h_{\alpha}(x)}=-\frac{s^{\prime}(x)}{s(\alpha)-s(x)}$ so that
$\mu^{\alpha}(x)$ $=\mu(x)-\dfrac{s^{\prime}(x)}{s(\alpha)-s(x)}\sigma^{2}(x),$ (3.6)
$\sigma^{\alpha}(x)$ $=\sigma(x),$
where we write $\mu^{\alpha}$ and $\sigma^{\alpha}$ instead of $\mu^{h_{\alpha}}$ and $\sigma^{h_{\alpha}}$ for simplicity.
Before proceeding further, we should check if the point $\alpha$ is an entrance boundary for the post-last passage time process $(X_{t})_{t>\lambda_{\alpha}}$: the process can enter into $(-\infty,\alpha)$ from $\alpha$ but cannot exit from $\alpha$. This confirms that $(X_{t})_{t>\lambda_{\alpha}}$ cannot return to the region $\lambda_{\alpha},\infty)$.
Proposition 3.3.
The point $\alpha$ is an entrance boundary for the post-last passage time process $X^{h_{\alpha}}_{t}(\omega):=\{\omega(t):\lambda_{\alpha}<t<H_{\ell},\mathbb{P}^{h_{\alpha}}_{\alpha}\}$ with $H_{\ell}:=\inf\{t:\omega(t)=\ell\}$.
Proof.
See Appendix A.4. ∎
4. Model Implementation
Now we apply the result in Section 3.2 to our asset-to-debt ratio process $Y$ and its post-last passage time process $Z$. Using (3.6) and the scale function $s(y)=\frac{1}{2M}(1-e^{-2M\cdot y})$ of the Brownian motion with drift $M$, we see that the process $Z_{t}$ follows the dynamics (4.1) right below; that is, the post-$\lambda_{\alpha^{*}}$ dynamics of the normalized asset-to-debt ratio process:
${\rm d}Z_{t}=M\coth\left(M\left(Z_{t}-\alpha^{*}\right)\right){\rm d}t+{\rm d}W_{t},\qquad Z_{0}=\alpha^{*}:=\frac{1}{\sigma}\ln(\alpha),\quad t>0.$ (4.1)
This is the process we shall work with for explicitly calculating LGD distribution. The starting point of $Z$ is $\alpha^{*}$. Note that since $\lim_{x\uparrow 0}\coth(x)=$-$\infty$, when the process $Z$ approaches $\alpha^{*}$ from below, the drift approaches negative infinity so that the process shall never reach the point $\alpha^{*}$ from the region $(-\infty,\alpha^{*})$. Mathematically, as proved in Proposition 3.3, $\alpha^{*}$ is an entrance boundary for $Z$ and financially, the firm’s asset-to-debt ratio process $Y$ shall never recover to the threshold level $\alpha$.
4.1. Loss-given-default distribution
The next task is to model the default time $\xi$. We define a random time that is suitable for default time. As is seen in the intensity-based default modeling, we need an exponential random variable $J$ with rate $\eta=1$, independent of everything else. We let $\Gamma:\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ be a nonnegative piecewise continuous function and define the moment inverse of integral functional by
$\nu(J):=\min\left(s:\int_{0}^{s}\Gamma\left(Z_{u}\right){\rm d}u=J\right).$ (4.2)
Following Bielecki and Rutkowski (2002, Sec.8.2),
Assumption 2.
We assume that $\Gamma$ satisfies
$\int_{0}^{\infty}\Gamma\left(Z_{s}\right){\rm d}s=\infty\quad\mathbb{P}-a.s.$ (4.3)
We can define random time $\tau:=\nu(J)$ and model the default time $\xi$ as
$\xi(\omega):=\lambda_{\alpha}(\omega)+\tau\circ\theta(\lambda_{\alpha}(\omega))$ (4.4)
where $\theta(\cdot)$ denotes the shift operator. Note that the construction of the random time is similar to the canonical construction of the default time in the intensity-based approach (see Proposition 5.26 in Capiński and Zastawniak (2017)). It follows from (4.2) that $\mathbb{P}\left(\tau=0\mid Z_{0}=\alpha^{*}\right)=0$ and from (4.3) that $\mathbb{P}\left(\tau<+\infty\mid Z_{0}=\alpha^{*}\right)=1$. We also see that $\mathbb{P}\left(\tau>s\mid Z_{0}=\alpha^{*}\right)>0$ for every $s\geq 0$. Moreover, we have $\mathbb{P}(\lambda_{\alpha}<\xi<+\infty)=1$ with (4.4). We have constructed the time $\tau$ using the process which never returns to $\alpha^{*}$, this feature being represented in (4.1).
We define LGD on debt $B$ as
$K^{B}(\xi):=1-Y_{\xi}=1-\frac{V_{\xi}}{B_{\xi}}.$ (4.5)
Due to $Y_{\xi}=Y_{\tau}\circ\theta(\lambda_{\alpha})$, it suffices to compute the distribution of $Z_{\tau}=\frac{1}{\sigma}\ln(Y_{\tau}\circ\theta(\lambda_{\alpha}))$. Thus, we are able to find the LGD distribution by focusing on $Z_{\tau}$ only. This simplifies the analysis as we do not need to consider the shift operator $\theta$ in the definition (4.4). Our model is summarized in Table 1: we have made only Assumptions 1 and 2.
Table 1. Comparison of our model with the conventional setting
Default time Distribution of the final value
Conventional $T=\inf\left\{t\geq 0:Y_{t}=c\right\}$ $Y_{T}=c$
Our approach $\xi=\lambda_{\alpha}+\tau\circ\theta({\lambda_{\alpha}})$ $Y_{\xi}\overset{d}{\sim}e^{\sigma Z_{\tau}}$
As shown in Proposition 3.3, $\alpha^{*}$ is an entrance boundary for $Z$. Hence we first consider an arbitrary starting position $z<\alpha^{*}$ and then take a limit. Fix any $Q<\alpha^{*}$. According to (4.1) and Theorem IV.5.1 in Borodin (2017),
$U(z):=\mathbb{E}_{z}\left[e^{-\gamma\tau}\mbox{1}\hskip-2.5pt\mbox{l}_{\{Z_{\tau}\leq Q\}}\right],\quad z<\alpha^{*},\quad\gamma>0$ (4.6)
is the unique continuous solution to the ordinary differential equation
$\frac{1}{2}U^{\prime\prime}(z)+M\coth(M(z-\alpha^{*}))U^{\prime}(z)-(\Gamma(z)+\gamma)U(z)=-\Gamma(z)\mbox{1}\hskip-2.5pt\mbox{l}_{[-\infty,Q]}(z),\quad z<\alpha^{*}.$ (4.7)
Here $\mathbb{E}_{z}[\cdot]$ denotes the conditional expectation $\mathbb{E}[\cdot\mid Z_{0}=z]$. We will provide an explicit form of $U(z)$ in Proposition 4.2. Using the solution $U(z)$, we obtain $\mathbb{E}_{\alpha^{*}}\left[e^{-\gamma\tau}\mbox{1}\hskip-2.5pt\mbox{l}_{\{Z_{\tau}\leq Q\}}\right]=\lim_{z\uparrow\alpha^{*}}U(z)$ by the continuity of $U(z)$. Finally, $\lim_{\gamma\downarrow 0}\mathbb{E}_{\alpha^{*}}\left[e^{-\gamma\tau}\mbox{1}\hskip-2.5pt\mbox{l}_{\{Z_{\tau}\leq Q\}}\right]$ produces the distribution of $K^{B}(\xi)$ via (4.5).
To obtain a concrete result, we specify the function $\Gamma$ in (4.2) as
$\Gamma\left(Z_{t}\right):=\mbox{1}\hskip-2.5pt\mbox{l}_{\{Z_{t}<\alpha^{*}\}}$ (4.8)
so that the integral in (4.2) represents the occupation time of the process $Z$ under the level $\alpha^{*}$. It follows that we identify the firm’s default once the asset-to-debt ratio process (in its logarithmic form) spends a certain random time below level $\alpha$ after time $\lambda_{\alpha}$. This specification is quite reasonable and natural since deterioration of credit quality is measured by the duration under water. As in other intensity-based models, the market observers are unaware of the firm’s state after a certain time (represented by $\lambda_{\alpha}$ in our model) and receive a default as a sudden shock.
Since the dynamics (2.5) is simple enough, we can compute the distribution of default time $\xi$ explicitly. From (4.2) and (4.8), we see that $\nu(J)=J$ since $Z$ does not hit the region $\alpha^{*},\infty)$: the occupation time of $(-\infty,\alpha^{*})$ after $\lambda_{\alpha}$ until $\xi$ is equal to $J$. This observation leads to the following result:
Proposition 4.1.
Under (4.8), the distribution of the default time $\xi$ is given by
$\mathbb{P}_{Y_{0}}(\xi\leq T)$ $=\int_{0}^{T}\left(\int_{0}^{T-x}\frac{p(u;\frac{1}{\sigma}\ln(Y_{0}),\alpha^{*})}{G(\alpha^{*},\alpha^{*})}{\rm d}u\right)e^{-x}{\rm d}x+\int_{0}^{T}\left(1-\frac{G(\frac{1}{\sigma}\ln(Y_{0}),\alpha^{*})}{G(\alpha^{*},\alpha^{*})}\right)e^{-x}{\rm d}x$ (4.9)
$=\frac{e^{-M(\frac{1}{\sigma}\ln(Y_{0})+\alpha^{*})}}{4|M|G_{0}(\alpha^{*},\alpha^{*})}e^{-|M(\frac{1}{\sigma}\ln(Y_{0})-\alpha^{*})|}\int_{0}^{T}e^{-x}\left(1+\mbox{Erf}\left[\frac{|M|(T-x)-|\frac{1}{\sigma}\ln(Y_{0})-\alpha^{*}|}{\sqrt{2(T-x)}}\right]-e^{2|M(\frac{1}{\sigma}\ln(Y_{0})-\alpha^{*})|}\mbox{Erfc}\left[\frac{|M|(T-x)+|\frac{1}{\sigma}\ln(Y_{0})-\alpha^{*}|}{\sqrt{2(T-x)}}\right]\right){\rm d}x$
$\hskip 28.45274pt+\left(1-\frac{G(\frac{1}{\sigma}\ln(Y_{0}),\alpha^{*})}{G(\alpha^{*},\alpha^{*})}\right)(1-e^{-T}),$
where
$\mbox{Erf}(x)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}e^{-t^{2}}{\rm d}t,\qquad\mbox{Erfc}(x)=1-\mbox{Erf}(x),$
and $G(\cdot,\cdot)$ is defined in (2.1).
Proof.
See Appendix A.5
As announced in the beginning of Section 2, we shall find $\alpha^{*}$ that matches (4.9) with $T=5$ with the market-quoted 5-year default probability.
Finally, we undertake the procedure to compute the distribution of the asset-to-debt ratio at time $\xi$ (i.e., $Y_{\xi}$), or equivalently that of its post-last passage time process at $\tau$ (i.e., $Z_{\tau}$) in view of (4.4).
Proposition 4.2.
Under (4.8), fix $\gamma>0$ and let $b_{1}=\sqrt{1+\frac{2(1+\gamma)}{M^{2}}}$. For $Q<\alpha^{*}$, we have
$\mathbb{E}_{\alpha^{*}}\left[e^{-\gamma\tau}\mbox{1}\hskip-2.5pt\mbox{l}_{\{Z_{\tau}\leq Q\}}\right]=\frac{1}{1+\gamma}\left(\cosh\left(|M|(\alpha^{*}-Q)\right)+b_{1}\sinh\left(|M|(\alpha^{*}-Q)\right)\right)e^{-b_{1}|M|(\alpha^{*}-Q)}.$ (4.10)
Proof.
The proofs of this proposition along with the subsequent corollaries are in Appendix A.6. ∎
Corollary 4.1.
The distribution of $Z_{\tau}$ is given by
$\mathbb{P}_{\alpha^{*}}\left(Z_{\tau}\leq Q\right)=\begin{cases}\left(\cosh\left(|M|(\alpha^{*}-Q)\right)+b_{2}\sinh\left(|M|(\alpha^{*}-Q)\right)\right)e^{-b_{2}|M|(\alpha^{*}-Q)},\quad Q\leq\alpha^{*}\\ \qquad\qquad\qquad\qquad\qquad 1,\qquad\qquad\qquad\qquad\qquad\qquad\qquad Q>\alpha^{*}\end{cases}$ (4.11)
where $b_{2}=\sqrt{1+\frac{2}{M^{2}}}$.
From this result we calculate the distribution of LGD $K^{B}(\xi)$ in (4.5):
Corollary 4.2.
The probability density function of $K^{B}(\xi)$ in (4.5) is given by
$\mathbb{P}_{\alpha^{*}}(K^{B}(\xi)\in{\rm d}x)=\frac{2}{|M|}\sinh\left(|M|\left(\alpha^{*}-\frac{\ln(1-x)}{\sigma}\right)\right)e^{-b_{2}|M|\left(\alpha^{*}-\frac{\ln(1-x)}{\sigma}\right)}\frac{1}{\sigma(1-x)}{\rm d}x,\quad 1-e^{\sigma\alpha^{*}}<x<1$
where $b_{2}=\sqrt{1+\frac{2}{M^{2}}}$. Moreover,
$\mathbb{E}_{\alpha^{*}}[K^{B}(\xi)]=1-\frac{e^{\sigma\alpha^{*}}}{1+\frac{\sigma^{2}}{2}+b_{2}\sigma|M|}.$ (4.12)
4.2. Parameter estimation
Fix any point in time as the current time. The simple procedure outlined below estimates the LGD distribution implied in the 5-year CDS spread using the information available up to the current time. The parameters are estimated in a way that the default probability and default time distribution obtained from the model are consistent with the default probability provided by LSEG Workspace (consistent with the quoted CDS spread).
  • Step 1:
    Estimate the parameters $(\mu,\sigma)$ of the asset process in (2.4). Since it is a well-known procedure, detailed explanation is delegated to Appendix A.7. The current value $Y_{0}$ of the asset-to-debt ratio process is retrieved using the estimated $\sigma$. The debt process $B$ represents the sum of short-term debt and a half of long-term debt (as in Moody’s KMV approach): the reason being due to the good fit to empirical evidence.
  • Step 2:
    Calibrate $\alpha$ so that the default time $\xi$ in (4.4) matches the market-quoted 5-year default probability by using Proposition 4.1.
  • Step 3:
    Obtain the LGD distribution of $K^{B}(\xi)$ from Corollary 4.2. However, in the case of CDS, it is more appropriate to consider the LGD distribution of the total debt $D=(D_{t})_{t\geq 0}$, which is done as follows:
    Letting the processes $(D^{S}_{t})_{t\geq 0}$ and $(D^{L}_{t})_{t\geq 0}$ represent short- and long-term debts respectively, we have $D_{t}=D^{S}_{t}+D^{L}_{t}$. For simplicity, we assume that the proportion of long-term debt $w=\frac{D_{t}^{L}}{D_{t}}$ remains constant. The loss rate of total debt $D$, denoted by $K^{D}(\xi)$, is then given by
    $K^{D}(\xi)=K^{B}(\xi)+\frac{1}{2}w(1-K^{B}(\xi)).$ (4.13)
4.3. Model Check
Finally, we make sure that the LGD distribution from our model is consistent with the CDS market. For this purpose, we shall calculate the 5-year CDS spread based on our model and compare it to the quoted spread. Recall that the underlying credit risk behind CDS spread refers to the risk of the whole company, not a particular debt obligation. (Therefore, this convention is consistent with our modeling.)
Note that the market CDS spread is calculated with the assumption of $60\%$ loss rate. In contrast, the value of the CDS spread we shall obtain below from our model depends on the LGD distribution. Therefore, the two CDS spreads cannot be compared directly. Instead we compare
$\rho:=\frac{\textnormal{CDS spread}}{\textnormal{average loss given default}}$ (4.14)
which is a spread per $1\%$ loss given default.
We set the principal amount to $\mathdollar 1$ and assume the spread payments are made quarterly. We find spread value for which the present value of spread payments (premium leg) and the present value of the payment in case of default (default leg) are equal. We ignore the counterparty risk and use the risk-free rate given by $r$ as the discount rate. The following procedure is for estimating $\rho$.
Let $\omega_{i}$, $i=1,\cdots,N$ denote a simulation path. Simulate a uniform random variable $U(\omega_{i})$ on $[0,1]$.
  • Step (i):
    We shall obtain samples of $K^{D}(\xi(\omega_{i}))$. First, simulate pairs $(\tau(\omega_{i}),Z_{\tau(\omega_{i})}(\omega_{i}))$. Recall that $\tau=J$ where $J$ is an exponential (with rate 1) random variable, so that we set $\tau(\omega_{i})=-\ln(U(\omega_{i}))$. By inverting the distribution function (4.11) at $U(\omega_{i})$, we obtain $Z_{\tau(\omega_{i})}(\omega_{i})$ which provides $K^{B}(\xi(\omega_{i}))=1-e^{\sigma Z_{\tau(\omega_{i})}(\omega_{i})}$ by (4.5) and, in turn, $K^{D}(\xi(\omega_{i}))$, by (4.13).
  • Step (ii):
    We shall obtain samples of $\xi(\omega_{i})$ by using $\tau(\omega_{i})$ in Step (i). Simulate $\lambda_{\alpha}(\omega_{i})$ by inverting its cumulative distribution function (which is also used in (4.9)) at $U^{\prime}(\omega_{i})$ (uniformly distributed random variable on $[0,1]$, independent of $U(\omega_{i})$). We thus obtain $\xi(\omega_{i})=\lambda_{\alpha}(\omega_{i})+\tau(\omega_{i})$.
  • Step (iii)
    We have obtained the pairs $(\xi(\omega_{i}),K^{D}(\xi(\omega_{i}))$ from which we can compute the present value of the recovery (on the dollar) upon default. Specifically, we calculate the value of $e^{-r\cdot\xi(\omega_{i})}K^{D}(\xi(\omega_{i}))\mbox{1}\hskip-2.5pt\mbox{l}_{\{\xi(\omega_{i})\leq 5\}}$. By taking the average over $N$ trials, we find the value of default leg.
  • Step (iv):
    Using the discount rate $r$, we calculate the present value of spread payments until $\xi(\omega_{i})$ (including accrued premium) or the maturity, whichever is earlier. By taking average over $N$ trials, we obtain the premium leg.
  • Step (v):
    We match the premium leg to the default leg to obtain the model-implied CDS spread.
  • Step (iv):
    Compute $\rho$ by (4.14) with $\mathbb{E}_{\alpha^{*}}\left[K^{D}(\xi)\mid\xi\leq 5\right]$ for its denominator. Compare this $\rho$ to the corresponding number from the CDS market to see these two $\rho$’s are close enough.
5. Empirical Study
5.1. Data
In this section, we analyze eight companies listed on New York Stock Exchange (NYSE). These firms together with their NYSE ticker symbols are as follows: Delta Air Lines, Inc. (DAL), Ford Motor Company (F), Kohl’s Corporation (KSS), Lennar Corporation (LEN), Toll Brothers, Inc. (TOL), Tyson Foods, Inc. (TSN), Whirlpool Corporation (WHR) and The Williams Companies, Inc. (WMB). In the sequel, we use NYSE ticker symbols for brevity. According to the Repository Reporting OTC Data on the Depository Trust and Clearing Corporation (DTCC) website, these firms have average daily notional amount of more than 15 million USD in CDS contracts during March 22, 2025 to June 20, 2025 (Q2)11endnote: 1https://www.dtcc.com/repository-otc-data (Accessed on 16/09/2025). Hence we believe that these names have sufficient liquidity in the CDS market. The selected companies are presented in Table 2 together with their industry classifications 22endnote: 2Industry classifications can be checked on Nasdaq website https://www.nasdaq.com/market-activity/stocks/screener (Accessed on 18/09/2025). The main purpose of presenting these examples is to illustrate how to implement our method to obtain the LGD distribution implied in the current credit market together with company-specific financial conditions. Specifically, we estimate the model parameters in a way that the model-implied default probability (which is based on default time distribution) is consistent with the default probability provided by LSEG Workspace. Based on those parameters, we obtain the LGD distribution. See Steps 1 through 3 in Section 4.2.
For each company, we choose one time point as the current day and use the information available up to that time point to obtain the LGD distribution implied in the market on that time point. We use 6-month daily data for the parameter estimation. The 6-month periods selected for each company are listed in Table 2. Since the purpose is to estimate the LGD distribution of relatively high credit risk scenarios, we selected the latest available 6-month periods which yield estimated drift parameter $M$ and barrier $\alpha$ satisfying $M<0$ and $\alpha\leq 1$.
We use the market yield on U.S. Treasury Securities at 5-year constant maturity (quoted on an investment basis, not seasonally adjusted) as the constant rate $r$. These rates were obtained from the Federal Reserve Bank of St. Louis website. We interpolated daily debt values using standardized quarterly balance sheets. We defined the debt $B$ as the sum of the short-term debt (notes payables, short-term debt, current portion of long-term debt and capital leases) and one half of the long-term debt (long-term debt and capital lease obligations). The average value of the ratio of long-term debt to total debt is given by $\widehat{w}$.
5.2. Procedure
Table 2 reports the estimated parameters of the asset-to-debt ratio processes; $\sigma,\mu,M$, and implied current $Y_{0}$ up to the first eleven rows. This is done by Step 1 in Section 4.2. The level of $\alpha$ is calibrated to match model-implied 5-year probability of default (PD) with the market counterpart (obtained from LSEG Workspace). See Step 2 of Section 4.2. The reader can see that the numbers in the line headed by ‘Market PD (5Y)’ are exactly the same as those in the line headed by ‘Model-implied PD (5Y)’ across the board: there is no arbitrariness in choosing $\alpha$. The third step is to compute $K^{D}(\xi)$ by Corollary 4.1. Figure 1 displays the density function of $K^{D}(\xi)$ in (4.13).
As a final procedure, we check the consistency of the model-implied LGD distribution with the current CDS market. Refer to Table 2 again. Following the procedure in Section 4.3 and setting the number of trials $N$ to 10,000, we calculate the current model-implied 5-year CDS spread based on the model-implied LGD distribution of $K^{D}(\xi)$. See the line headed by ‘Estimated CDS spread (5Y)’. Note that in the calculation process we use samples of $K^{D}(\xi(\omega_{i}))$ and $\xi(\omega_{i})$ from Step (i) and Step (ii) of Section 4.3, respectively. For each company, we calculate model-implied $\rho$ (‘Model $\rho$’) and compare it to the market $\rho$ (‘Market $\rho$’) which is the ratio of quoted CDS spread over 60% loss rate.
Figure 2 shows the density function of $\lambda_{\alpha^{*}}$, where $\alpha^{*}=\frac{1}{\sigma}\ln(\alpha)$. Recall the normalized asset-to-debt ratio process $\frac{1}{\sigma}\ln(Y)$ in (2.5) with state space $(\ell,r)=(-\infty,+\infty)$ whose killing boundary, based on which the $\lambda_{\alpha^{*}}$ density function is computed, is the left natural boundary $\ell=$-$\infty$; no killing in the interior. In Figure 3, we provide the density of the occupancy time above $\alpha^{*}$ before killing, which is denoted by $\lambda_{\alpha^{*}}^{A,c}$. The process $\frac{1}{\sigma}\ln(Y)$ is forcibly killed on the first instant to level $c^{*}=\frac{1}{\sigma}\ln(c)$. The density of $\lambda_{\alpha^{*}}^{A,c}$ shows, based on the firm-value approach, how much time the asset-to-debt ratio stays above $\alpha^{*}$ before it hits $c^{*}$. From the viewpoint of risk management, we can set $c^{*}$ at a landmark point for decision making. In this analysis, we set $c=0$, $\frac{2}{3}\alpha$, and almost equal to $\alpha$. The case $c\rightarrow 0$ corresponds to $c^{*}\rightarrow$-$\infty$, which is one of the natural boundaries of $\frac{1}{\sigma}\ln(Y)$. By comparing $\lambda_{\alpha^{*}}$ and $\lambda_{\alpha^{*}}^{A,c=0}$, we see how much time the process $\frac{1}{\sigma}\ln(Y)$ spends above level $\alpha^{*}=\frac{1}{\sigma}\ln(\alpha)$, relative to the last passage time $\lambda_{\alpha^{*}}$. See the end of this section (as well as A.7) for the computation of Figures 2 and 3.
Table 2. Estimated parameters up to 4 decimal points together with the standard error of $\ln(\sigma)$ produced by the maximum likelihood estimation. Company industry classifications in order: Meat/Poultry/Fish, Homebuilding, Natural Gas Distribution, Homebuilding, Air Freight/Delivery Services, Consumer Electronics/Appliances, Auto Manufacturing, Department/Specialty Retail Stores.
TSN TOL WMB LEN DAL WHR F KSS
Dataset start date 2024-12-11 2025-01-29 2022-11-29 2024-11-25 2024-12-24 2024-12-24 2024-12-04 2025-01-30
Dataset end date=Current Date 2025-06-16 2025-07-31 2023-06-01 2025-05-30 2025-06-30 2025-06-30 2025-06-09 2025-08-01
Current risk-free rate 0.0404 0.0396 0.037 0.0396 0.0379 0.0379 0.0409 0.0377
$\hat{w}$ 0.9317 1.0000 0.9374 1.0000 0.8376 0.7135 0.6496 0.8399
Estimated $\ln(\sigma)$ -1.6378 -1.0399 -1.6882 -1.1320 -0.8134 -1.3957 -2.3723 -0.9104
Standard Error 0.0630 0.0631 0.0630 0.0630 0.0631 0.0640 0.0241 0.1003
Log-Likelihood -903.7338 -885.8554 -980.2036 -1005.2365 -1072.2375 -812.9502 -1029.2806 -711.4242
$\sigma$ 0.1944 0.3535 0.1849 0.3224 0.4433 0.2477 0.0933 0.4024
$\mu$ -0.1728 -0.2310 -0.1741 -0.9825 -0.2969 -0.0741 -0.0117 -0.3108
$M$ -1.1937 -0.9423 -1.2344 -3.3316 -0.9769 -0.5761 -0.6106 -1.0672
Implied Current $Y_{0}$ 5.1854 8.7477 3.9325 18.0141 4.7479 2.1385 1.3728 1.3409
$\alpha$ 0.9967 0.8307 0.9215 0.0592 0.2689 0.8190 0.9598 0.3541
Expected value of $K^{D}(\xi)$ 0.6138 0.7503 0.6411 0.9867 0.9160 0.6261 0.4355 0.8855
Market PD (5Y) 4.04 6.824 8.177 8.445 9.062 13.95 16.208 58.142
Model-implied PD (5Y) 4.0400 6.8241 8.1769 8.4451 9.0621 13.9499 16.2077 58.1419
Market CDS spread (5Y) 47.29 83.14 96.05 99.52 136.68 171.51 201.4 925.32
Estimated CDS spread (5Y) 41.4160 88.6270 90.3355 152.7450 153.7701 159.6580 130.6061 1182.6249
5Y PD from CDS estimation (Step(v)) 4.04 6.82 8.17 8.44 9.06 13.96 16.2 58.16
Market $\rho$ 0.7882 1.3857 1.6008 1.6587 2.2780 2.8585 3.3567 15.4220
Model $\rho$ 0.7520 1.2829 1.5425 1.5617 1.7220 2.7382 3.1847 13.5751
(Model $\rho$ - Market $\rho$)/Market $\rho$ -0.0459 -0.0742 -0.0365 -0.0584 -0.2441 -0.0421 -0.0512 -0.1198
5.3. Analysis
The market data is consistent with the fact that higher default probability corresponds to higher CDS spread. Estimated CDS spread is based on the distribution of $K^{D}(\xi)$ which is different from constant 60% loss rate used as market convention; therefore, estimated and market spreads will naturally differ. However, as shown in the bottom row of Table 2, for six out of eight companies their percentage difference of Model and Market $\rho$’s is less than $7.42\%$. We use Table 2 together with Figure 1 for the comparison. Figure 1 suggests that for ‘TSN’, ‘WMB’, and ‘WHR’, $K^{D}(\xi)$ is distributed somewhat around 60%. These are three companies whose expected loss rate (‘Expected value of $K^{D}(\xi)$’) is closest to 60%. We look at the bottom row of Table 2 to find that these three companies have the smallest percentage difference (in absolute value) between ‘Market $\rho$’ and ‘Model $\rho$’, no more than 4.6%. Based on these observations, we can state that the model is consistent with the market. The exceptions are ‘DAL’ and ‘KSS’. For these two companies, the deviation of ‘Market $\rho$’ and ‘Model $\rho$’ is the largest (24% and 12% respectively). In the case of very high CDS spread levels (like ‘KSS’), prices are more susceptible to liquidity (supply-demand gap) than to fundamental credit quality; and hence prices become volatile and unpredictable. Hence the deviation of ‘KSS’ ‘Market $\rho$’ from ‘Model $\rho$’ is understandable. For the remaining case of ‘DAL’, it seems reasonable to attribute the discrepancy of the $\rho$’s to the high volatility $\sigma=0.4433$, which may contribute to the difficulty in pricing spreads. Moreover, we point out an observation common to ‘DAL’ and ‘KSS’: the distribution of their $K^{D}(\xi)$ is also concentrated at much higher levels (around 90%) than 60%. From these, we derive a conclusion that the closer the distribution of $K^{D}(\xi)$ is to 60% loss rate, the estimated spread per 1% loss rate (which is the definition of $\rho$) is sufficiently close to the market counterpart. On the other hand, this conclusion suggests that for companies whose loss rate distribution is concentrated in a region far from 60%, pricing based on the distribution of $K^{D}(\xi)$ will provide significantly different results and this point should be carefully taken into consideration. Let us take a close look at $\lambda_{\alpha^{*}}$ (Figure 2) and $\lambda^{A,c}_{\alpha^{*}}$ (Figure 3) of the two exceptional ‘DAL’ and ‘KSS’. The density of both $\lambda_{\alpha^{*}}$ and $\lambda_{\alpha^{*}}^{A,c}$ in the case of ‘DAL’ is concentrated around 6 years, around the same levels of the other (six) companies while ‘KSS’ has concentration at much lower level. Taking into account that the $K^{D}(\xi)$ of these companies have similar concentration, $K^{D}(\xi)$ and $\lambda_{\alpha^{*}}$ (or $\lambda_{\alpha^{*}}^{A,c}$) provide different aspects of credit risk management.
Figure 1. Probability density function of total debt LGD $K^{D}(\xi)$ in (4.13) based on estimated parameters
Figure 2. Probability density function ${\rm d}\mathbb{P}_{y^{*}}(\lambda_{\alpha^{*}}\leq t)/{\rm d}t$, with $y^{*}=\frac{1}{\sigma}\ln(Y_{0})$, of the last passage time $\lambda_{\alpha^{*}}$ based on estimated parameters (including $\alpha$ and the starting point $Y_{0}$)
Looking at Figure 3 again, we see that the shape of the density for ‘LEN’ and ‘KSS’ is different from that of the remaining companies (all of which have a similar shape). The same pattern in observed in Figure 2. For these two companies, the densities for both $\lambda_{\alpha^{*}}$ and $\lambda_{\alpha^{*}}^{A,c}$ are concentrated either within much narrower range (‘LEN’) or at much lower level (‘KSS’), compared to other companies, indicating high probability of the asset-to-debt ratio process not remaining above $\alpha^{*}$ for much longer. Based on Table 2, ‘LEN’ has the lowest drift $M$ and the highest expected value of $K^{D}(\xi)$ while ‘KSS’ has the highest 5Y PD. Thus, Figures 2 and 3 accurately capture high credit risk. Let us compare Figures 2 and 3 for all the companies; more specifically, Figure 2 and the density with $c=0$ in Figure 3. We see that densities of $\lambda_{\alpha^{*}}$ and $\lambda_{\alpha^{*}}^{A,c}$ with $c=0$ put relatively similar masses in the same regions, which means that the process $\frac{1}{\sigma}\ln(Y)$ spends most of the time above level $\alpha^{*}$ up until the last passage time $\lambda_{\alpha^{*}}$. This is due to the higher initial values $Y_{0}$ as compared to $\alpha$, especially for the companies with low CDS spreads. But it is equally true that once the process goes below $\alpha^{*}$, the chance it recovers to $\alpha^{*}$ is small. The last observation can be confirmed by the densities with $c=0$ (in blue) and $c=\frac{2\alpha}{3}$ (in orange): they are almost identical.
Finally, in drawing the graphs in this section, we used the explicit formula in Corollary 4.2 and (4.13) for Figure 1. For Figure 2, we used, in writing the normalized version (with *), $\mathbb{P}_{x^{*}}(\lambda_{\alpha^{*}}\in{\rm d}t)=\frac{p\left(t;x^{*},\alpha^{*}\right)}{s(\alpha^{*})-s(\ell)}{\rm d}t$ where $p(t;x^{*},z^{*})=\frac{1}{2\sqrt{2\pi t}}\exp\left(-M(x^{*}+z^{*})-\frac{M^{2}t}{2}-\frac{(x^{*}-z^{*})^{2}}{2t}\right)$, $s(x^{*})=\frac{1}{2M}(1-e^{-2Mx^{*}})$ and $\ell=$-$\infty$. The initial value $x^{*}$ in the above equation is $\frac{1}{\sigma}\ln(Y_{0})$. In Figure 3, we consider the occupation time of the level above $\alpha^{*}$ before the process $\frac{1}{\sigma}\ln(Y)$ hits some level $\frac{1}{\sigma}\ln(c)=c^{*}<\alpha^{*}$. We denote this time by $\lambda_{\alpha^{*}}^{A,c}$. Its Laplace transform can be easily calculated using (3.1) and is given $\mathbb{E}_{x^{*}}\left[e^{-q\lambda_{\alpha^{*}}^{A,c}}\right]=\frac{\frac{e^{-(\sqrt{M^{2}+2q}+M)x^{*}}}{k}}{\frac{e^{-(\sqrt{M^{2}+2q}+M)\alpha^{*}}}{k}+\frac{1}{2M}(e^{-2Mc^{*}}-e^{-2M\alpha^{*}})}$ with $k=(\sqrt{M^{2}+2q}+M)e^{-(\sqrt{M^{2}+2q}-M)\alpha^{*}}$ as derived in Appendix A.2.
Figure 3. Probability density function (produced by inverting Laplace transform) of the occupancy time above $\alpha^{*}$ before the normalized asset-to-debt ratio hits level $c^{*}$ (starting at $y^{*}=\frac{1}{\sigma}\ln(Y_{0})$) based on the estimated parameters. Three levels of $c$ (stated as legends) are considered for each company: 0, $\frac{2}{3}\alpha$, and $\alpha-0.001$.
Appendix A
A.1. Excursion-theoretic (Point process) view
We can view $\{X_{t},t>\lambda_{\alpha}\}$ as an excursion from $\alpha$ of infinite length. To pursue this point of view, we review some of the excursion theory based on Getoor (1979) and Rogers and Williams (1994b), in particular its Sections VI. 50, 54 and 55.
Let us return to the original process $X$ and consider an excursion from a regular point $\alpha$. Let the local time at $y\in\mathcal{I}$ be $L(t,y),t\geq 0$. Set $L(t)=L(t,\alpha)$ and define the right inverse $\rho(t,\alpha):=\inf\{s:L(s)>t\}$. It is well known that $\rho$ is a right-continuous and nondecreasing subordinator under $\mathbb{P}_{\alpha}$ and hence
$\mathbb{E}_{\alpha}\left[e^{-q\rho(t,\alpha)}\right]=e^{-tg(q)}$ (A.1)
where
$g(q)=m\{\alpha\}q+\int_{(0,\infty]}(1-e^{-qs})\nu({\rm d}s)$ (A.2)
with $m\{\alpha\}=0$ in our case and $\nu$ being a measure on $(0,\infty]$ satisfying $\int(1\wedge s)\nu({\rm d}s)<\infty$.
By using the tail distribution
$c(u):=\nu((u,\infty]),$
the Laplace exponent is rewritten as
$g(q)=m\{\alpha\}q+q\int_{0}^{\infty}e^{-qs}c(s){\rm d}s.$ (A.3)
Essentially, $\rho(t,\alpha)$ corresponds to the length of excursion from point $\alpha$ that occurs at local time $t$, so that $\lim_{u\rightarrow\infty}c(u)=\nu(\{+\infty\})$ corresponds to an excursion of infinite length.
We define
$r_{q}(x,y):=\mathbb{E}_{x}\left(\int_{0}^{\infty}e^{-qt}{\rm d}L(t,y)\right),\quad x,y\in\mathcal{I},q>0.$
Then, by a change of variable and (A.1),
$r_{q}(\alpha,\alpha)=\mathbb{E}_{\alpha}\left(\int_{0}^{\infty}e^{-qt}{\rm d}L(t,y)\right)=\mathbb{E}_{\alpha}\left(\int_{0}^{\infty}e^{-q\rho(t,\alpha)}{\rm d}t\right)=\frac{1}{g(q)}.$ (A.4)
But if we define the $q$-potential $U^{q}f(x):=\mathbb{E}_{x}\left(\int_{0}^{\infty}e^{-qt}f(X_{t}){\rm d}t\right)$ of a bounded function $f$ with $x\in\mathcal{I}$, as is written in the proof of Theorem V.50.7 in Rogers and Williams (1994b),
$U^{q}f(x)=\int_{\mathcal{I}}m({\rm d}y)f(y)\mathbb{E}_{x}\left(\int_{0}^{\infty}e^{-qs}{\rm d}L(s,y)\right)$
which is also written as
$U^{q}f(x)=\int_{\mathcal{I}}m({\rm d}y)f(y)G_{q}(x,y)$
by using $G_{q}(x,y)=\int_{0}^{\infty}e^{-qt}p(t;x,y){\rm d}t$. See (2.3) for its exact form. By comparison, $G_{q}(x,y)=\mathbb{E}_{x}\left(\int_{0}^{\infty}e^{-qs}{\rm d}L(s,y)\right)$ a.e. $y$. But by the right continuity in $y$, we have the equality for every $y\in\mathcal{I}$. We obtain from (A.4) that $g(q)=\frac{1}{G_{q}(\alpha,\alpha)}$ and hence
$\mathbb{E}_{\alpha}(e^{-q\rho(t,\alpha)})=\exp\left(-\frac{t}{G_{q}(\alpha,\alpha)}\right).$
Remark A.1.
Let $X$ be a linear diffusion with regular point $\alpha\in\mathcal{I}$. We consider the local time at $\alpha$ and excursions from this point. With $g(q)=\frac{1}{G_{q}(\alpha,\alpha)}$, (A.3) becomes
$\int_{0}^{\infty}e^{-qs}c(s){\rm d}s=\frac{1}{qG_{q}(\alpha,\alpha)},$
which can be used to obtain $c(u)=\nu((u,\infty])$ through the inverse Laplace transform. This $c(u)$ corresponds to the rate at which the first excursion of length greater than $u$ occurs. To consider an excursion of infinite length, which is closely related to a post-last passage time process discussed in Section 3.2, let us take $q\downarrow 0$ in (A.1) to find
$\mathbb{P}_{\alpha}(\rho(t,\alpha)<\infty)=e^{-t\nu(\{+\infty\})}$
by (A.2). Moreover, from $g(q)=\frac{1}{G_{q}(\alpha,\alpha)}$, we have
$\nu(\{+\infty\})=\lim_{q\downarrow 0}\frac{1}{G_{q}(\alpha,\alpha)}:=\frac{1}{G_{0}(\alpha,\alpha)}.$
$\diamondsuit$
A.2. On computations when non-killing (interior) state is made to be killing
Let us comment on the killing boundary involved in the computation of (3.1). Take an example of (I) Brownian motion with negative drift $M$ (as is the normalized $Y$ process in Section 2.3) with two natural boundaries $\ell=$-$\infty$ and $r=+\infty$, of which $\ell$ is a killing boundary. We, of course, use the Green functions $G(\alpha,\alpha)$ and $G_{q}^{A}(x,\alpha)$ for this process to compute the Laplace transform (3.1). Next, let us consider the case (II) where we set a killing boundary, other than natural boundaries, at some $c\in(\ell,r)=\mathcal{I}$. Then, we should use the q-Green function $G_{q}^{A}(x,\alpha)$ for the process “Brownian motion with drift on $(c,+\infty)$ killed at cc”, which is not the same as that for simple Brownian motion with drift.
For case (I), the increasing and decreasing solutions to $\mathcal{G}f=qf$ are given by $\psi_{q}(x)=e^{(\sqrt{M^{2}+2q}-M)x}$ and $\phi_{q}(x)=e^{-(\sqrt{M^{2}+2q}+M)x}$. For case (II), let these solutions be denoted by $\psi_{q}^{c}$ and $\phi_{q}^{c}$ which are linear combinations of $\psi_{q}$ and $\phi_{q}$. For the condition at the killing boundary $c$ we refer the reader to Borodin and Salminen (2002, Chapter II.1. Setion 10). This condition is given by
$\psi_{q}^{c}(c+)=0.$
We have $\psi_{q}^{c}(c)=a_{1}\psi_{q}(c)+a_{2}\phi_{q}(c)$ from which we obtain
$a_{1}=e^{-(\sqrt{M^{2}+2q}+M)c},\quad a_{2}=-e^{(\sqrt{M^{2}+2q}-M)c}.$
As there is no boundary condition for $\phi_{q}^{c}$, we set $\phi_{q}^{c}(x)=\phi_{q}(x)$. Thus, we have
$\psi^{c}_{q}(x)=a_{1}\psi_{q}(x)+a_{2}\phi_{q}(x)=e^{-(\sqrt{M^{2}+2q}+M)c}e^{(\sqrt{M^{2}+2q}-M)x}-e^{(\sqrt{M^{2}+2q}-M)c}e^{-(\sqrt{M^{2}+2q}+M)x}$
$\phi_{q}^{c}(x)=e^{-(\sqrt{M^{2}+2q}+M)x}$
$(\phi_{q}^{c})^{+}(x)=-(\sqrt{M^{2}+2q}+M)e^{-(\sqrt{M^{2}+2q}-M)x}$
together with Wronskian $w_{q}^{c}=2\sqrt{M^{2}+2q}\cdot e^{-(\sqrt{M^{2}+2q}+M)c}$. Then, $G_{q}^{A}(x,\alpha)$ with $x\geq\alpha$ for case (II) is given by
$\frac{-1}{(\phi_{q}^{c})^{+}(\alpha)}\phi_{q}^{c}(x)\left(\frac{-(\phi_{q}^{c})^{+}(\alpha)}{w_{q}^{c}}\psi_{q}^{c}(\alpha)+\frac{(\psi_{q}^{c})^{+}(\alpha)}{w_{q}^{c}}\phi_{q}^{c}(\alpha)\right)=\frac{e^{(\sqrt{M^{2}+2q}-M)\alpha}\cdot e^{-(\sqrt{M^{2}+2q}+M)x}}{(\sqrt{M^{2}+2q}+M)}.$
Finally, as the introduction of the killing boundary $c$ does not change the scale function, case (II) Green function is given by $G^{c}(\alpha,\alpha)=s(\alpha)-s(c)$. This is equal to the Green function of the process in case (I) with $\ell$ just being replaced by $c$. This equality is in fact true for every diffusion, not just Brownian motion with drift. For the details of $G^{A}_{q}(\cdot,\cdot)$, we refer the reader to Egami and Kevkhishvili (2025).
A.3. Proof of Lemma 3.1
Define $B(x)=\int^{x}\frac{2}{\sigma^{2}(y)}\mu(y){\rm d}y$. It is well-known that
$m(x)=\frac{2}{\sigma^{2}(x)}e^{B(x)}\quad\text{and}\quad s^{\prime}(x)=e^{-B(x)}$
and
$m^{h}({\rm d}y)=h^{2}(y)m({\rm d}y)\quad\text{and}\quad s^{h}({\rm d}y)=\frac{1}{h^{2}(y)}s({\rm d}y),$
where $m^{h}(\cdot)$ and $\sigma^{h}(\cdot)$ are the speed measure and scale function of the $h$-transform of $X$. See Section II.1 of Borodin and Salminen (2002). By simple comparison, we obtain
$\frac{1}{(\sigma^{h})^{2}(x)}e^{B^{h}(x)}=h^{2}(x)\frac{1}{\sigma^{2}(x)}e^{B(x)}\quad\text{and}\quad e^{-B^{h}(x)}=\frac{1}{h^{2}(x)}e^{-B(x)}$ (A.5)
where $B^{h}(x)=\int^{x}\frac{2}{(\sigma^{h})^{2}(y)}\mu^{h}(y){\rm d}y$. From (A.5), we have $\sigma^{h}(x)=\sigma(x)$. This in turn gives
$\exp\left(2\int^{x}\frac{\mu^{h}(y)-\mu(y)}{\sigma^{2}(y)}{\rm d}y\right)=h^{2}(x),$
so that $\mu(x)$ satisfies $\mu^{h}(x)=\mu(x)+\frac{h^{\prime}(x)}{h(x)}\sigma^{2}(x)$.
A.4. Proof of Proposition 3.3
For the post-last passage time process, we need only to consider a point $y<\alpha$. In view of (3.5), (3.3) becomes
$p^{h_{\alpha}}(t;x,y)=\frac{s(\alpha)-s(y)}{s(\alpha)-s(x)}p(t;x,y).$
This means that the post-last passage time process is obtained equivalently via the transform $g(\cdot):=s(\alpha)-s(\cdot)$. Call this transform also $X^{h_{\alpha}}$ and its scale function and speed measure $s^{g}$ and $m^{g}$, respectively. By the general result for the transform of this kind (see Appendix A.3), we have
$m^{g}({\rm d}y)=g^{2}(y)m({\rm d}y)\quad\text{and}\quad(s^{g})^{\prime}(y)=\frac{1}{g^{2}(y)}s^{\prime}(y),$
which gives us $s^{g}(y)=\frac{1}{s(\alpha)-s(y)}$. Then, $s^{g}(\alpha)-s^{y}(y)=\infty$ for $y<\alpha$ and
$\int_{y}^{\alpha}(s^{g}(\eta)-s^{g}(y))m^{g}({\rm d}\eta)$ $=\int_{y}^{\alpha}\left(\frac{1}{s(\alpha)-s(\eta)}-\frac{1}{s(\alpha)-s(y)}\right)(s(\alpha)-s(\eta))^{2}m({\rm d}\eta)$
$<\int_{y}^{\alpha}\frac{1}{s(\alpha)-s(\eta)}(s(\alpha)-s(\eta))^{2}m({\rm d}\eta)$
$=\int_{y}^{\alpha}(s(\alpha)-s(\eta))m({\rm d}\eta)<\infty.$
The last finiteness result is due to $\alpha$ being a regular point for $X$. This proves that $\alpha$ is an entrance boundary for the $g$-transform, which is equivalent to the post-last passage time process $X^{h_{\alpha}}$ (see [Karlin and Taylor (1981), Chap.15, Table 6.2]).
A.5. Proof of Proposition 4.1
Using the fact that $J$ is distributed exponentially with rate 1 and independent of $\lambda_{\alpha}$, we have for $T\in\mathbb{R}_{+}$
$\mathbb{P}_{Y_{0}}(\xi\leq T)$ $=\mathbb{P}_{Y_{0}}(\lambda_{\alpha}\leq T-J)=\int_{0}^{T}\mathbb{P}_{Y_{0}}(\lambda_{\alpha}\leq T-x)e^{-x}{\rm d}x$
$=\int_{0}^{T}\mathbb{P}_{Y_{0}}(0<\lambda_{\alpha}\leq T-x)e^{-x}{\rm d}x+\int_{0}^{T}\mathbb{P}_{Y_{0}}(\lambda_{\alpha}=0)e^{-x}{\rm d}x.$
From (2.5) and (4.1), $\lambda_{\alpha}=\sup\{t:\frac{1}{\sigma}\ln(Y_{t})=\alpha^{*}\}$. The last passage time distribution is known: Proposition 4 in Salminen (1984) and Proposition 3.1 in Egami and Kevkhishvili (2020) provide
$\mathbb{P}_{y}(\lambda_{\alpha}\in{\rm d}u)=\frac{p(u;y,\alpha)}{G(\alpha,\alpha)}{\rm d}u\quad\text{and}\quad\mathbb{P}_{y}(\lambda_{\alpha}=0)=1-\frac{G(y,\alpha)}{G(\alpha,\alpha)}$
where $G(\alpha,\alpha)$ is available in (2.1). It can be confirmed that $\mathbb{P}_{Y_{0}}(\lambda_{\alpha}=0)=0$ when $Y_{0}\geq\alpha$ by (2.1) and Assumption 1. Since the transition density $p$ for the process $\frac{1}{\sigma}\ln(Y_{t})$, a Brownian motion with drift $M$, is available, we can evaluate the integral in the above equation explicitly to obtain (4.9).
A.6. Proof of Proposition 4.2
Define $Z^{*}_{t}:=\alpha^{*}-Z_{t}$. Then, by (4.1)
${\rm d}Z^{*}_{t}=-{\rm d}Z_{t}=-M\coth(M(Z_{t}-\alpha^{*})){\rm d}t-{\rm d}W_{t}=M\coth(MZ^{*}_{t}){\rm d}t+{\rm d}W_{t}^{*}=|M|\coth(|M|Z^{*}_{t}){\rm d}t+{\rm d}W_{t}^{*}$
where $(W^{*}_{t})_{t\geq 0}:=(-W_{t})_{t\geq 0}$ is a standard Brownian motion. We have used the fact that $M\coth(-M\cdot x)=M\coth(M\cdot x)$. We have $\tau=\min\left(s:\int_{0}^{s}\mbox{1}\hskip-2.5pt\mbox{l}_{\{Z_{u}<\alpha^{*}\}}{\rm d}u=J\right)=\min\left(s:\int_{0}^{s}\mbox{1}\hskip-2.5pt\mbox{l}_{\{Z_{u}^{*}>0\}}{\rm d}u=J\right)$ and
$\mathbb{E}_{\alpha^{*}}\left[e^{-\gamma\tau}\mbox{1}\hskip-2.5pt\mbox{l}_{\{Z_{\tau}\leq Q\}}\right]$ $=\lim_{z^{\prime}\uparrow\alpha^{*}}\mathbb{E}\left[e^{-\gamma\min\left(s:\int_{0}^{s}\mbox{1}\hskip-2.04861pt\mbox{l}_{\{Z_{u}<\alpha^{*}\}}{\rm d}u=J\right)}\mbox{1}\hskip-2.5pt\mbox{l}_{\{Z_{\tau}\leq Q\}}\mid Z_{0}=z^{\prime}\right]$
$=\lim_{z\downarrow 0}\mathbb{E}\left[e^{-\gamma\min\left(s:\int_{0}^{s}\mbox{1}\hskip-2.04861pt\mbox{l}_{\{Z_{u}^{*}>0\}}{\rm d}u=J\right)}\mbox{1}\hskip-2.5pt\mbox{l}_{\{Z^{*}_{\tau}\geq\alpha^{*}-Q\}}\mid Z_{0}^{*}=z\right].$ (A.6)
By this translation to $Z^{*}$, (4.6) becomes $U(z)=\mathbb{E}\left[e^{-\gamma\min\left(s:\int_{0}^{s}\mbox{1}\hskip-2.04861pt\mbox{l}_{\{Z_{u}^{*}>0\}}{\rm d}u=J\right)}\mbox{1}\hskip-2.5pt\mbox{l}_{\{Z^{*}_{\tau}\geq\alpha^{*}-Q\}}\mid Z_{0}^{*}=z\right]$ and (4.7) is now
$\frac{1}{2}U^{\prime\prime}(z)+|M|\coth(|M|\cdot z)U^{\prime}(z)-\left(1+\gamma\right)U(z)$ $=-\mbox{1}\hskip-2.5pt\mbox{l}_{[\alpha^{*}-Q,+\infty)}(z),\quad z\in(0,\infty).$ (A.7)
(i) Consider the case when $z\geq\alpha^{*}-Q$. Then (A.7) is reduced to
$\frac{1}{2}U^{\prime\prime}(z)+|M|\coth(|M|\cdot z)U^{\prime}(z)-\left(1+\gamma\right)U(z)=-1.$
According to Section IV.16.7 and Appendix 2.12 in Borodin (2017), its fundamental solutions are given by
$\psi_{1}(z)=\frac{\sqrt{2}\sinh\left(|M|b_{1}z\right)}{\sqrt{\pi}b_{1}\sinh(|M|\cdot z)},\quad\phi_{1}(z)=\frac{\sqrt{\pi}e^{-|M|b_{1}z}}{\sqrt{2}\sinh(|M|\cdot z)},$ (A.8)
where $b_{1}=\sqrt{1+\frac{2(1+\gamma)}{M^{2}}}$. Here $\psi_{1}$ is an increasing solution and it satisfies $\lim_{z\downarrow 0}\psi_{1}(z)=\sqrt{\frac{2}{\pi}}$ and $\lim_{z\uparrow+\infty}\psi_{1}(z)=+\infty$ because $b_{1}>1$. On the other hand, $\phi_{1}$ is a decreasing solution and it satisfies $\lim_{z\downarrow 0}\phi_{1}(z)=+\infty$ and $\lim_{z\uparrow+\infty}\phi_{1}(z)=0$. Therefore, for $z\geq\alpha^{*}-Q$, we have $U(z)=\frac{1}{1+\gamma}+c\phi_{1}(x)$ with some constant $c$. We have eliminated the first fundamental solution because $U(z)$ needs to be bounded when $z\to+\infty$.
(ii) Let $0<z<\alpha^{*}-Q$. In this case, (A.7) is reduced to
$\frac{1}{2}U^{\prime\prime}(z)+|M|\coth(|M|\cdot z)U^{\prime}(z)-(1+\gamma)U(z)=0.$
As in the case (i) above, its fundamental solutions are given by $\psi_{1}(z)$ and $\phi_{1}(z)$ in (A.8). Therefore, in this case we have $U(z)=d\psi_{1}(z)$ with some constant $d$. We have eliminated the second fundamental solution because $U(z)$ is bounded when $z\to 0$.
Finally, from the continuity of $U(z)$ and $U^{\prime}(z)$ at $\alpha^{*}-Q$, we determine the constants $c$ and $d$. In particular,
$d=\frac{1}{1+\gamma}\sqrt{\frac{\pi}{2}}\left(\cosh\left(|M|(\alpha^{*}-Q)\right)+b_{1}\sinh\left(|M|(\alpha^{*}-Q)\right)\right)e^{-b_{1}|M|(\alpha^{*}-Q)}.$
Finally, $\lim_{z\downarrow 0}U(z)=\lim_{z\downarrow 0}d\cdot\psi_{1}(z)=d\cdot\sqrt{\frac{2}{\pi}}$, which proves (4.10) in view of (A.6).
A.6.1. Proof of Corollary 4.1
Note that $\mathbb{P}_{\alpha^{*}}\left(Z_{\tau}\leq Q\right)=1$ for $Q>\alpha^{*}$ due to (4.1). For $Q\leq\alpha^{*}$, we have
$\mathbb{P}_{\alpha^{*}}\left(Z_{\tau}\leq Q\right)=\lim_{\gamma\downarrow 0}\mathbb{E}_{\alpha^{*}}\left[e^{-\gamma\tau}\mbox{1}\hskip-2.5pt\mbox{l}_{\{Z_{\tau}\leq Q\}}\right].$
Thus, we obtain the desired result by taking the limit $\gamma\downarrow 0$ in (4.10).
A.6.2. Proof of Corollary 4.2
Due to $Y_{\xi}\overset{d}{\sim}e^{\sigma Z_{\tau}}$, we have $K^{B}(\xi)\overset{d}{\sim}1-e^{\sigma Z_{\tau}}$ by (4.5). The density function of $Z_{\tau}$ is obtained by differentiating (4.11) with respect to $Q$:
$f(Q):=\frac{\mathbb{P}_{\alpha^{*}}(Z_{\tau}\in{\rm d}Q)}{{\rm d}Q}=\frac{2}{|M|}\sinh(|M|(\alpha^{*}-Q))e^{-b_{2}|M|(\alpha^{*}-Q)}{\rm d}Q,\quad Q<\alpha^{*}.$
Then, the density of the transform $1-e^{\sigma Z_{\tau}}$ is given by
$\mathbb{P}_{\alpha^{*}}(K^{B}(\xi)\in{\rm d}x)=f\left(\frac{\ln(1-x)}{\sigma}\right)\frac{1}{\sigma(1-x)}{\rm d}x,\quad 1-e^{\sigma\alpha^{*}}<x<1.$
Finally, $\int_{1-e^{\sigma\alpha^{*}}}^{1}x\mathbb{P}_{\alpha^{*}}(K^{B}(\xi)\in{\rm d}x)$ provides (4.12).
A.7. Detailed comments on Step 1 in Section 4.2
We estimate the parameters of the asset process in (2.5) using the data of the firm’s equity $E$ and debt $B$ available up to the current time. The debt process $B$ represents a certain amount of debt to be repaid and we let it be the sum of short-term debt and a half of long-term debt (as in Moody’s KMV approach) since empirical evidence has demonstrated that such level is an appropriate default threshold for the firm’s assets.
We use the option-theoretic approach of Merton (1974) where the value of equity represents the value of the European call option written on the firm’s assets with the strike price equal to the future value of debt. We estimate the asset parameters using the maximum likelihood method of Duan (1994), Duan (2000), and Lehar (2005). We assume that debt $B$ grows at the risk-free rate $r$. Assuming that debt $B$ is exogenous, the log-likelihood function to be maximized is given by
$L(E_{1},\cdots,E_{n};\mu,\sigma\mid B_{1},\cdots B_{n})=-\frac{n-1}{2}\ln(2\pi)-\frac{n-1}{2}\ln(\sigma^{2}\Delta_{t})-\sum_{t=2}^{n}\ln(\hat{V}_{t})-\sum_{t=2}^{n}\ln(\Phi(\hat{d}_{t}))$
$-\frac{1}{2\sigma^{2}\Delta_{t}}\sum_{t=2}^{n}\left(\ln\left(\frac{\hat{V}_{t}}{\hat{V}_{t-1}}\right)-\left(\mu-\frac{1}{2}\sigma^{2}\right)\Delta_{t}\right)^{2}$
where $\Phi$ is a standard normal cumulative distribution function and $\Delta_{t}$ is the time interval between two consecutive data points. We use $n$ number of past data points for the estimation. In the above equation, $\hat{V}_{t}$ is the solution to
$E_{t}=V_{t}\Phi(d_{t}(V_{t}))-B_{t}\Phi(d_{t}(V_{t})-\sigma\sqrt{T_{m}})$
with respect to $V_{t}$ and $\hat{d}_{t}=d_{t}(\hat{V}_{t})$ where
$d_{t}(V_{t})=\frac{\ln\left(\frac{V_{t}}{B_{t}}\right)+\frac{\sigma^{2}}{2}T_{m}}{\sigma\sqrt{T_{m}}}.$
As in Lehar (2005), we set the maturity $T_{m}$ of the call option to 1 year. By differentiating the log-likelihood function with respect to $\mu$ and setting this partial derivative to zero, we obtain
$\mu=\frac{\sum_{t=2}^{n}\ln\left(\frac{\hat{V}_{t}}{\hat{V}_{t-1}}\right)}{(n-1)\Delta_{t}}+\frac{1}{2}\sigma^{2}.$
We may plug this into the log-likelihood function which then becomes a function of only one parameter $\sigma$. By maximizing the log-likelihood, we obtain the estimate of $\sigma$ and as a byproduct, the estimate of $\mu$ as well.
Acknowledgment
We thank Tomohiro Koike (Graduate School of Economics, Kyoto University) for his valuable comments regarding Proposition 3.2.